### v2

# Rydberg superatoms: An artificial quantum system for quantum information processing and quantum optics

Abstract
Dense atom ensembles with Rydberg excitations display intriguing collective effects mediated by their strong, long-range dipole-dipole interactions. These collective effects, often modeled using Rydberg superatoms, have gained significant attention across various fields due to their potential applications in quantum information processing and quantum optics. In this review article, we delve into the theoretical foundations of Rydberg interactions and explore experimental techniques for their manipulation and detection. We also discuss the latest advancements in harnessing Rydberg collective effects for quantum computation and optical quantum technologies. By synthesizing insights from theoretical studies and experimental demonstrations, we aim to provide a comprehensive overview of this rapidly evolving field and its potential impact on the future of quantum technologies.
I INTRODUCTION
Rydberg atoms exhibit striking properties that arise from their large atomic size and strong interatomic interactions.Gallagher (1988, 1994); Saffman, Walker, and Mølmer (2010) The huge induced electric dipole moment couples with external electric fields, allowing the manipulation of the electronic as well as motional states of Rydberg atoms with laser and microwave (MW) fields. Importantly, strong and long-range dipole-dipole or van der Waals (vdW) interactions are found between Rydberg atoms separated by macroscopic distances of over 10 micrometers.Urban et al. (2009); Gaëtan et al. (2009) Such interactions can be controlled by laser or MW fields. This paves the way for realizing scalable quantum information processing architectures, Jaksch et al. (2000) creating novel many-body states Browaeys and Lahaye (2020) and exploring nonlinear quantum optics.Firstenberg, Adams, and Hofferberth (2016)
At the heart of the Rydberg physics lies strong interaction-induced collective behaviors.Heidemann et al. (2007); Dudin et al. (2012); Zeiher et al. (2015) In high-lying Rydberg states, the interatomic interaction (1111100100100100 MHz) can be much stronger than typical laser Rabi frequencies (tens of kHz–a few MHz).Saffman, Walker, and Mølmer (2010) As a result, the laser excitation of one atom inhibits the excitation of neighboring atoms within a certain distance (i.e. blockade radius). Within the blockade radius, only the many-body ground state and singly excited Rydberg state can be occupied. The former consists of atoms in the electronic ground state, while the latter is a collective state where one atom is in the Rydberg state while all other atoms are in the electronic ground state. Dynamics of such collective states are conveniently described by an artificial superatom.Lukin et al. (2001) This description can be extended to many-atom settings, where the collective effect allows for the creation of exotic many-body correlated states.Pohl, Demler, and Lukin (2010); Olmos, González-Férez, and Lesanovsky (2009); Honer et al. (2010); Weimer et al. (2010); Ji, Ates, and Lesanovsky (2011); de Léséleuc et al. (2019) These potentials make Rydberg atoms a leading candidate for implementing quantum simulation and quantum computing tasks. Moreover, strong Rydberg atom interactions can be mapped to optical fields, generating sizable optical nonlinearities.Pritchard et al. (2010) These have found wide usage in exploring nonlinear quantum optics Peyronel et al. (2012); Firstenberg, Adams, and Hofferberth (2016) and realizing optical quantum devices, such as single-photon sources,Dudin and Kuzmich (2012); Ripka et al. (2018); Padrón-Brito et al. (2021); Ornelas-Huerta et al. (2020); Shi et al. (2022) photonic quantum logic gates,Tiarks et al. (2016, 2019); Stolz et al. (2022) photonic entanglement filter,Ye et al. (2023) single-photon absorber,Tresp et al. (2016) single-photon switch,Baur et al. (2014) and single-photon transistors.Gorniaczyk et al. (2014)
In recent years, quite a few review papers have meticulously surveyed the rapidly evolving field of Rydberg atom technologies,Saffman, Walker, and Mølmer (2010); Browaeys and Lahaye (2020); Saffman (2016); Schauss (2018); Adams, Pritchard, and Shaffer (2020); Beterov et al. (2020); Henriet et al. (2020); Morgado and Whitlock (2021); Wu et al. (2021); Shi (2022); Yuan et al. (2023a) emphasizing their diverse applications and groundbreaking advancements in quantum computation, quantum simulation, and quantum sensing, showcasing the versatility and potential of Rydberg atoms. This review article distinguishes itself by concentrating specifically on the concept of Rydberg superatoms, with a particular emphasis on their role as a quantum interface between atoms and photons. Rydberg superatoms are created through collective excitation states, where multiple atoms collectively display characteristics akin to a single two-level atom. This collective behavior not only enhances the coherence and controllability of quantum states but also endows the interface with novel electromagnetic properties, enabling the creation and manipulation of photon-photon interactions at the quantum level. This review thus will contribute to the existing body of knowledge on Rydberg research, offering valuable insights and guidance for future studies.
In this review article, we delve into the theory underpinning the Rydberg superatom, related collective effects, and key experimental achievements. In particular, we review the harnessing of Rydberg collectivity in two critical areas: quantum information processing and quantum optics. We start with an introduction to Rydberg atoms and Rydberg superatoms in Sec. II, discussing the atomic properties, two-body interactions of Rydberg atoms, and the emergence of collective states described by Rydberg superatoms. In Sec. III, coherent laser excitation of Rydberg atoms and the preparation of Rydberg superatoms are covered, including quantum interference of Rydberg superatoms. In Sec. IV, we discuss quantum information processing based on Rydberg superatoms, focusing on implementing quantum logic gates, generating quantum entanglement, and performing quantum simulations. We describe the application of Rydberg superatoms in quantum optics in Sec. V, including the demonstration and realization of strong nonlocal nonlinear optics, single-photon applications, and hybridizing quantum systems with Rydberg superatoms. Finally, we conclude in Sec. VI.
II RYDBERG ATOMS AND RYDBERG SUPERATOMS
II.1 Characteristics of Rydberg atoms
In a broad sense, a Rydberg atom can be defined as an atom with one of its valence electrons in a highly excited state characterized by a large principal quantum number. The category of “atoms” that exhibit this characteristic includes alkali atoms, Gallagher (1988); Saffman, Walker, and Mølmer (2010); Lim, Lee, and Ahn (2013) alkaline earth atoms,Nakagawa, Kawakami, and Ueda (2018); Cooper et al. (2018); Zhang et al. (2020a); Schäfer et al. (2020); Kaufman and Ni (2021); Norcia and Ferlaino (2021); Mukherjee et al. (2011); Wilson et al. (2022); Piñeiro Orioli, Thompson, and Rey (2022); Muni et al. (2022); Wu et al. (2022) trapped ions,Müller et al. (2008); Feldker et al. (2015); Li and Lesanovsky (2012); Higgins et al. (2017); Gambetta, Lesanovsky, and Li (2019); Zhang et al. (2020b); Gambetta et al. (2021) as well as solid state systems.Kazimierczuk et al. (2014); Aßmann and Bayer (2020); Heckötter et al. (2017) Since alkali atoms possess only one valence electron, their properties in the Rydberg state closely resemble those of a hydrogen atom. This allows us to derive the distinctive properties of alkali Rydberg atoms using, for example, the quantum defect theory. Gallagher (1988, 1994) The binding energy of the Rydberg states of alkali atoms (e.g., Rb and Cs) with the principal quantum number $n$, orbital angular momentum $l$, and total angular momentum $j=l\pm 1/2$ can be expressed as
$E_{n,l,j}=-\frac{hcR_{m}}{(n-\delta_{n,l,j})^{2}}.$ (1)
Here, $R_{m}=R_{\infty}/(1+m_{e}/m_{\rm{nucleus}})$ represents the mass-corrected Rydberg constant, where $m_{e}$ denotes the electron mass and $m_{\rm{nucleus}}$ refers to the atomic mass of the nucleus. The term $R_{\infty}=m_{e}e^{4}/(8\epsilon_{0}^{2}h^{3}c)$ is the Rydberg constant, where $e$ is the elementary charge, $\epsilon_{0}$ denotes the permittivity of the vacuum, $h$ represents the Planck constant, $c$ stands for the speed of light in vacuum.Seaton (1983); Gallagher (1994) Furthermore, the species-dependent quantum defect, denoted as $\delta_{n,l,j}$, can be approximated by the modified Rydberg-Ritz parameters
$\delta_{n,l,j}=\delta_{0}+\frac{\delta_{2}}{(n-\delta_{0})^{2}}+\frac{\delta_{4}}{(n-\delta_{0})^{4}}+\frac{\delta_{6}}{(n-\delta_{0})^{6}}+...,$ (2)
where the coefficients $\delta_{0}$, $\delta_{2}$, $\delta_{4}$, and $\delta_{6}$ are primarily depends on $l$ (while also being related to $j$) and remains non-negligible solely for states with $l\leq 3$.Lorenzen and Niemax (1983); Gallagher (1994) Precise understanding of the quantum defect is paramount for anticipating spectral lines, comprehending electron behavior in highly excited states, and probing the boundaries of quantum mechanics. This imperative has led researchers to pursue quantum defect parameters of Rydberg atoms with high precision through various methodologies.Goy et al. (1982); Lorenzen and Niemax (1984); Weber and Sansonetti (1987); Gallagher (1988); Dyubko et al. (1995); Li et al. (2003); Han et al. (2006); Afrousheh et al. (2006); Sanguinetti et al. (2009); Mack et al. (2011); Deiglmayr et al. (2016); Sautenkov et al. (2016); Li et al. (2019a, 2021); Bai et al. (2023) In Table 1, we present a comprehensive compilation of modified Rydberg-Ritz parameters for low orbit states in Rb and Cs atoms. This compilation incorporates data from early seminal work and recent measurements. For additional orbitals or other alkali atoms, their quantum defects can be easily obtained using two recently released open-source software packages, pairinteractionWeber et al. (2017) and Alkali Rydberg Calculator (ARC) toolbox.Šibalić et al. (2017)
Table 1: Quantum defect constants $\delta_{0}$, $\delta_{2}$, $\delta_{4}$, and $\delta_{6}$ for different orbital angular momentum $l$ of Rb and Cs extracted from Refs. Lorenzen and Niemax, 1983, 1984; Li et al., 2003; Deiglmayr et al., 2016; Li et al., 2021; Bai et al., 2023.
Rydberg series $\delta_{0}$ $\delta_{2}$ $\delta_{4}$ $\delta_{6}$
85Rb $nS_{1/2}$ 3.1311804Li et al. (2003) 0.1784Li et al. (2003) -1.8Lorenzen and Niemax (1983)
85Rb $nP_{3/2}$ 2.64142Li et al. (2021) 0.295Li et al. (2021) 0.97495Lorenzen and Niemax (1983) 14.6001Lorenzen and Niemax (1983)
85Rb $nD_{5/2}$ 1.3464572Li et al. (2003) -0.596Li et al. (2003) -1.50517Lorenzen and Niemax (1983) -2.4206Lorenzen and Niemax (1983)
85Rb $nF_{7/2}$ 0.016411Li et al. (2021) -0.0784Li et al. (2021) -0.36005Lorenzen and Niemax (1983) 3.2390Lorenzen and Niemax (1983)
133Cs $nS_{1/2}$ 4.0493532Deiglmayr et al. (2016) 0.23915Deiglmayr et al. (2016) 0.06Deiglmayr et al. (2016) 11Deiglmayr et al. (2016)
133Cs $nP_{3/2}$ 3.5590676Deiglmayr et al. (2016) 0.37469Deiglmayr et al. (2016) -0.67431Lorenzen and Niemax (1984) 22.3531Lorenzen and Niemax (1984)
133Cs $nD_{5/2}$ 2.4663144Deiglmayr et al. (2016) 0.01381Deiglmayr et al. (2016) -0.392Deiglmayr et al. (2016) -1.9Deiglmayr et al. (2016)
133Cs $nF_{7/2}$ 0.0335646Bai et al. (2023) -0.2052Bai et al. (2023)
Starting from the expression for Rydberg’s binding energy, we can obtain a series of scaling laws that exhibit outstanding characteristics concerning the principal quantum number $n$, or more precisely, the effective principal quantum number $n^{*}=(n-\delta_{n,l,j})$. The energy spacing between adjacent Rydberg levels is given by
$\Delta E=E_{n+1,l+1,j}-E_{n,l,j}\approx\frac{2hcR_{m}[(\delta_{n,l,j}-\delta_{n,l+1,j})+1]}{n^{*3}},$ (3)
which follows a $n^{*-3}$ scaling. The corresponding transition wavelength typically falls within the millimeter-wave band between 10 mm (30 GHz) and 1 mm (300 GHz).Šibalić et al. (2017) This range of wavelengths offers exceptional precision in spectroscopic measurements, Prozument et al. (2011); Gordon et al. (2014) showing transient amplifiers of microwave radiation Raimond et al. (1982); Moi et al. (1983); Goy et al. (1983) and contributing to the development of the fifth generation mobile communication network.Deb and Kjærgaard (2018); Yuan et al. (2023b) It serves a crucial role in the advancement of atomic physics and quantum science research.Carter and Martin (2013); Kiffner et al. (2016); Grimes et al. (2017); Wade et al. (2017); Peper et al. (2019); Suleymanzade et al. (2020); Kanungo et al. (2022); Legaie, Raithel, and Anderson (2023)
Figure 1: (a) The numerical data demonstrates the scaling laws for the average orbit radius of Rydberg atoms in the $S$ orbit (blue circles) and circular orbit (red squares), expressed in terms of the effective principal quantum number denoted as $n^{*}$ for each case. (b) The scaling laws govern the radial dipole matrix elements for transitions between neighboring Rydberg states, denoted as $|\langle nP_{3/2}|e\hat{r}|nS_{1/2}\rangle|$ (blue diamonds), as well as for transitions between a low-lying state and a Rydberg state, represented as $|\langle nP_{3/2}|e\hat{r}|5S_{1/2}\rangle|$ (red hexagons), where $n^{*}=n-\delta_{n,2,j}$ for the $P$ orbit.
The expectation value of the electron orbit, determined by the radial wave function for an alkali atom state with medium to high quantum numbers $n$, is given byLindgård and Nielsen (1977); Theodosiou (1984)
$\langle\hat{r}\rangle_{n^{*},l,j}=\frac{1}{2}[3n^{*2}-l(l+1)]a_{0},$ (4)
where $a_{0}\approx 5.29\times 10^{-2}~{}\rm~{}nm$ represents the Bohr radius. This formula offers an exact description of the radius of Rydberg atomic orbitals with high principal quantum numbers, as the Coulomb potential predominates in this regime. The Rydberg atomic wave function is influenced by quantum defects. They mainly generate a phase shift to the wave function. Hence wave functions of alkali atoms are different from that of hydrogen atoms as long as quantum defects are non-negligible. As a result, the expectation value of physical quantities can be directly derived from the hydrogen wave function, with the only alteration being the substitution of the principal quantum number $n$ with the effective principal quantum number $n^{*}$.Kazimierczuk et al. (2014); Heckötter et al. (2017) For low $l$ states, particularly $l=0$ ($S$ states), the expectation value simplifies to be $23n^{*2}a_{0}/$ 2, but for the circular orbit with $l=n-1$, where the quantum defect is zero, the value becomes $(n^{2}+n/2)a_{0}$. For large $n$ the radius of the Rydberg electron can be hundreds of nanometers. This permits the formation of ultra-long range Rydberg molecules, Herzberg (1987); Greene, Dickinson, and Sadeghpour (2000); Bendkowsky et al. (2009, 2010); Kiffner et al. (2012); Krupp et al. (2014); Camargo et al. (2016); Li et al. (2011); Eiles and Greene (2017); Eiles et al. (2017); Shaffer, Rittenhouse, and Sadeghpour (2018); Whalen et al. (2020); Giannakeas et al. (2020); Peper and Deiglmayr (2021); Christian Fey and Schmelcher (2020); Hummel, Eiles, and Schmelcher (2021) where ground state atoms are trapped within the Rydberg electron cloud. In Fig. 1(a), the ARC toolbox Šibalić et al. (2017) is used to perform numerical calculation of the expectation values for the $S$ and circular orbits of Rubidium, encompassing the principal quantum number $n$ from 10 to 100. The data exhibit quadratic scaling with $n^{*}$, closely aligning with the analytical expression.
[…]
Lifetimes of Rydberg atoms are much longer than typical time scales of laser coupling and Rydberg atom interactions. This is important for maintaining quantum coherence, a crucial factor for quantum computation and simulation. The spontaneous decay rate from $|n,l,j\rangle$ to lower energy state $|n^{\prime},l^{\prime},j^{\prime}\rangle$ is governed by the Einstein coefficient (in SI units):
$A_{nlj\rightarrow n^{\prime}l^{\prime}j^{\prime}}$ $=$ $\frac{2\omega^{3}_{n^{\prime}l^{\prime}j^{\prime},nlj}}{3\epsilon_{0}hc^{3}}(2j^{\prime}+1)l_{\rm max}\left\{\begin{array}[]{ccc}l,&l^{\prime},&1\\ j^{\prime},&j,&1/2\end{array}\right\}^{2}$ (8)
×$\times|\langle n^{\prime},l^{\prime},j^{\prime}|e\hat{r}|{n},l,j\rangle|^{2},$
where $\omega_{n^{\prime}l^{\prime}j^{\prime},nlj}$ denotes the corresponding transition angular frequency, $l_{max}$ signifies the larger of the two values between $l$ and $l^{\prime}$, and the curly braces indicate a Wigner-$6j$ symbol.Seiler et al. (2016); Beterov et al. (2009) The radial part of Rydberg electronic wave functions has a vanishing overlap with that of ground and weakly excited states, decreasing $A_{nlj\rightarrow n^{\prime}l^{\prime}j^{\prime}}$. At zero temperature, the overall decay rate of Rydberg states to many low-lying states is proportional to $n^{*-3}$, where in this case $\omega_{n^{\prime}l^{\prime}j^{\prime},nlj}$ is nearly constant and only the electric dipole transitions dominate. However, in circular Rydberg states ($l=n-1$), because $A_{nlj\rightarrow n^{\prime}l^{\prime}j^{\prime}}\propto\omega^{3}_{n^{\prime}l^{\prime}j^{\prime},nlj}$, they can only undergo a transition to an adjacent Rydberg state via a dipole-allowed process.Hulet and Kleppner (1983); Nguyen et al. (2018); Hölzl et al. (2024) This leads to the spontaneous emission rate of the circular state being proportional to $n^{*-5}$.
Compared to the high $l$ states, the quantum defects of the low $l\leq 3$ states have a greater impact on the energy levels correction. Consequently, the weak external static electric field will not cause the Rydberg atom in the low orbit state to endure a direct transition between the adjacent orbits but will instead present the energy level shift as a second-order perturbation, which is in the form of a quadratic Stark effectKhadjavi, Lurio, and Happer (1968); Angel and Sandars (1968)
$\Delta E_{\rm Stark}=\sum_{n^{\prime},l^{\prime}j^{\prime}\neq n,l,j}\frac{|\langle n^{\prime},l^{\prime},j^{\prime},m_{j}^{\prime}|e\hat{r}\cdot\bm{E}|n,l,j,m_{j}\rangle|^{2}}{E_{n^{\prime},l^{\prime},j^{\prime}}-E_{n,l,j}}=-\frac{1}{2}\alpha_{0}E^{2}$ (9)
where we have assumed the static electric field $\bm{E}$ is applied along the $z$-axis creating off-diagonal couplings between states, with the selection rule $\Delta l=\pm 1$, $\Delta m_{j}=0$. As determined by the Wigner-Eckart theorem,Weissbluth (1978) the dipole matrix elements are linked to the radial matrix elements $\langle n^{\prime},l^{\prime},j^{\prime}|e\hat{r}|{n},l,j\rangle$ through coefficients independent of the principal quantum number. Therefore, the scalar polarizability $\alpha_{0}$ scales with $n^{*7}$ due to the combined scaling of the radial dipole matrix element and the energy spacing for neighboring Rydberg states in the denominator.Zimmerman et al. (1979); O’Sullivan and Stoicheff (1985); Lai et al. (2018); Bai et al. (2020a); Chakraborty and Sahoo (2023) It should be noted that states possessing quantum numbers of angular momentum $l>3$ manifest degeneracy, leading to a linear Stark shift. For a more precise description of the Stark structure of the Rydberg states in alkali atoms, direct diagonalization of the energy matrix is required. The Rydberg atoms due to high polarizability can be exploited as highly accurate sensors for electric fieldsSedlacek et al. (2012); Fan et al. (2014, 2015); Anderson, Paradis, and Raithel (2018); Cox et al. (2018); Meyer et al. (2020); Jing et al. (2020); Crawford et al. (2021); Meyer et al. (2021); Calero et al. (2021); Hu et al. (2022); Liu et al. (2022); Fancher et al. (2023); Liu et al. (2023a); Ding et al. (2022); Yuan et al. (2023a) and the nondestructive imaging of atoms and ions.Günter et al. (2012); Gross, Vogt, and Li (2020)
Table 2: Scaling of various properties of Rydberg atoms with the effective principal quantum numbers $n^{*}$. The data for the transition rate in the parentheses is scaling in the circular state.
Property Expression $n^{*}$ scaling
Binding energy $E_{n,l,j}=-\frac{{hcR_{m}}}{(n-\delta_{n,l,j})^{2}}$ $n^{*-2}$
Level spacing $E_{n+1,l+1,j}-E_{n,l,j}$ $n^{*-3}$
Orbit radius $\langle\hat{r}\rangle_{n^{*},l,j}=\frac{1}{2}[3n^{*2}-l(l+1)]a_{0}$ $n^{*2}$
Dipole moment $\langle nP_{3/2}|e\hat{r}|n(5)S_{1/2}\rangle$ $n^{*2}$ ($n^{*-1.5}$)
Transition rate $A_{nlj\rightarrow n^{\prime}l^{\prime}j^{\prime}}$ $n^{*-3}(n^{*-5})$
Scalar polarizability $\alpha_{0}$ $n^{*7}$
In Table 2, a concise overview is presented, detailing the scaling behavior of various properties in Rydberg atoms with respect to the effective principal quantum number. A comprehensive exploration of additional Rydberg atom properties and associated mechanisms can be found in numerous recent review articles and references therein.Saffman, Walker, and Mølmer (2010); Löw et al. (2012); Browaeys, Barredo, and Lahaye (2016); Saffman (2016); Schauss (2018); Wüster and Rost (2018); Adams, Pritchard, and Shaffer (2020); Beterov et al. (2020); Browaeys and Lahaye (2020); Henriet et al. (2020); Morgado and Whitlock (2021); Wu et al. (2021); Shi (2022)
II.2 Interaction potential between Rydberg atoms
The interaction potential between Rydberg atoms manifests as an electrostatic interaction, which can be expressed as an infinite series of powers involving the inverse of the interatomic distance $R$.Dalgarno and Davison (1966) The predominant component is represented by the lowest-order dipole termsJaksch et al. (2000); You and Chapman (2000)
$\hat{V}_{int}=\frac{1}{4\pi\epsilon_{0}}[\frac{\hat{{d_{1}}}\cdot\hat{{d_{2}}}-3(\hat{{d_{1}}}\cdot\bm{n})(\hat{{d_{2}}}\cdot\bm{n})}{R^{3}}],$ (10)
where $\hat{d_{i}}=-e\hat{r_{i}}$ is the electric dipole operator of atom $i$ ($i=1,2$), and $\bm{n}$ denotes the unit vector connecting the two atoms. Without loss of generality, the polar angle between the quantization axis $z$ and $\bm{n}$ is set to $\theta$, and the azimuthal angle is $\phi$. In the spherical basis, the electric dipole-dipole interaction can be expressed as
$\hat{V}_{int}$ $=$ $\frac{1}{4\pi\epsilon_{0}R^{3}}[\frac{1-3\cos^{2}\theta}{2}(\hat{d}_{1}^{+}\hat{d}_{2}^{-}+\hat{d}_{1}^{-}\hat{d}_{2}^{+}+2\hat{d}_{1}^{0}\hat{d}_{2}^{0})$ (11)
$+\frac{3}{\sqrt{2}}\sin\theta\cos\theta(\hat{d}_{1}^{+}\hat{d}_{2}^{0}e^{i\phi}-\hat{d}_{1}^{-}\hat{d}_{2}^{0}e^{-i\phi}+\hat{d}_{1}^{0}\hat{d}_{2}^{+}e^{i\phi}$
$-\hat{d}_{1}^{0}\hat{d}_{2}^{-}e^{-i\phi})-\frac{3}{2}\sin^{2}\theta(\hat{d}_{1}^{+}\hat{d}_{2}^{+}e^{2i\phi}+\hat{d}_{1}^{-}\hat{d}_{2}^{-}e^{-2i\phi})],$
where $\hat{d}^{0}=\hat{d}^{z}$, $\hat{d}^{+}=-(\hat{d}^{x}+i\hat{d}^{y})/\sqrt{2}$, and $\hat{d}^{-}=(\hat{d}^{x}-i\hat{d}^{y})/\sqrt{2}$, which correspond to changes in magnetic quantum number of $\Delta m_{j}=0$, $+1$, and $-1$, respectively, in the process of atomic transition. This formula enables us to construct typical interaction forms for the Rydberg atom system. In the following, we will discuss different interaction types encountered in Rydberg atom experiments.
[…]
$\hat{H}_{\rm spin}=\frac{(1-3\cos^{2}\theta)}{2}\frac{C_{3}}{R^{3}}(|\uparrow\downarrow\rangle\langle\downarrow\uparrow|+{\rm H.c.}),$ (12)
where ${\rm H.c.}$ stands for the Hermitian conjugate and the real coefficient $C_{3}=\langle\uparrow\downarrow\left|\hat{d}_{1}^{+}\hat{d}_{2}^{-}|\downarrow\uparrow\right\rangle/(4\pi\epsilon_{0})$. This Hamiltonian was initially employed to examine the coherent dynamics of a spin excitation in a chain comprising three Rydberg atoms with $\theta=0$. Barredo et al. (2015); de Léséleuc et al. (2017) Subsequently, it was extended to simulate a topological model protected by symmetry in a zigzag configuration of Rydberg atoms, with the magic angle $\theta=\arccos(1/\sqrt{3})$ governing the hopping direction along the next nearest neighbor atoms. de Léséleuc et al. (2019) Additionally, it was utilized to synthesize a density-dependent Peierls phase employing three Rydberg atoms arranged in an equilateral configuration with $\theta=\pi/2$. Lienhard et al. (2020) Furthermore, this model finds application in the simulation of energy transportSchönleber et al. (2015) and various generic spin chain models.Geier et al. (2021); Scholl et al. (2022); Chen et al. (2023a); Nishad et al. (2023); Chen et al. (2023b); Bornet et al. (2024); Li, Li, and Shao (2024)
When the electronic energy of a pair of Rydberg atoms is (nearly) degenerate with energies of other Rydberg pair states, the dipole-dipole interaction can be strong due to the so-called Föster resonance. Even if a small energy defect exists, it can be mitigated by the differential Stark effect induced by external electric fields. Walker and Saffman (2005); Ryabtsev et al. (2010) The mathematical expression closely resembles Eq. (12), differing only in the composition of the bases. Experimental observations indicate that in the presence of a modest electric field ($F_{\rm res}\approx 33$ mV/cm), resonance occurs between the state $|dd\rangle$ and the symmetric state $|pf_{1}\rangle_{\rm s}=(|pf_{1}\rangle+|f_{1}p\rangle)/\sqrt{2}$ of two 87 Rb atoms, Ravets et al. (2014, 2015) where $|d\rangle=|59D_{3/2},m_{j}=3/2\rangle$, $|p\rangle=|61P_{1/2},m_{j}=1/2\rangle$ and $|f_{1}\rangle=|57F_{5/2},m_{j}=5/2\rangle$. Therefore, this interaction channel is effectively isolated from other interaction channels at an interatomic distance $R=9.1~{}\mu\rm m$ in the presence of a 3.3 G magnetic field. The corresponding Hamiltonian is given by
$\hat{H}_{\rm F\ddot{o}rster}=\frac{(1-3\cos^{2}\theta)}{\sqrt{2}}\frac{C_{3}}{R^{3}}(|dd\rangle_{\rm s}\langle pf_{1}|+{\rm H.c.}),$ (13)
with the measured $C_{3}/h=2.39\pm 0.03~{}{\rm GHz}~{}{\mu}{\rm m}^{3}$. A more systematic study of Förster resonance involving different atomic species can be found in Ref. Beterov and Saffman, 2015, and the Stark-tuned three-body Förster resonances have been predicted Faoro et al. (2015) and observed.Tretyakov et al. (2017); Ryabtsev et al. (2018) By integrating adiabatic rapid passage and Landau-Zener control with Förster resonance tuning, robust two-qubit quantum logic gates can be achieved. Beterov et al. (2016a, 2018); Huang et al. (2018) Additionally, a recent experimental breakthrough has demonstrated an ultrafast Förster oscillation capable of performing a conditional phase gate on nanosecond timescales. Chew et al. (2022)
In the absence of an external electric field, the energy difference defined by $\Delta_{r^{\prime},r^{\prime\prime}}=E_{r^{\prime}}+E_{r^{\prime\prime}}-2E_{r}$ of the Rydberg atom pair in the Förster resonance process greatly exceeds the strength of the dipole-dipole interaction when the atoms are at a significant distance from each other (perturbation theory is applicable). Here, $E_{r}$ represents the energy of the Rydberg state $|r\rangle$ of interest, while $E_{r^{\prime}}$ and $E_{r^{\prime\prime}}$ denote the energies of a pair of Rydberg states $|r^{\prime}\rangle$ and $|r^{\prime\prime}\rangle$, whose symmetric state is dipole coupled to the Rydberg pair state $|rr\rangle$. This leads to a multichannel-induced Stark shift in the energy levels of the atom pair, resulting in the interaction between the atoms manifesting as a classical vdw interaction,Walker and Saffman (2008); Saffman, Walker, and Mølmer (2010) represented by
$\hat{H}_{\rm vdw}=-\frac{C_{6}}{R^{6}}|rr\rangle\langle rr|.$ (14)
Here, $C_{6}$ is defined as $C_{6}=\sum_{r^{\prime}r^{\prime\prime}}|\langle r^{\prime}r^{\prime\prime}|\hat{V}_{int}|rr\rangle|^{2}/{\Delta_{r^{\prime},r^{\prime\prime}}}$, where in principle $|r^{\prime}r^{\prime\prime}\rangle$ encompasses all pair states dipole coupled to $|rr\rangle$. However, for practical computational efficiency, focusing on quantum states within a specific range of energy differences allows one to obtain sufficiently accurate values of $C_{6}$. When the interatomic distance is not significantly large and perturbation theory is not applicable, it becomes crucial to perform the diagonalization of the Hamiltonian governing the pair-state interaction. Experimentally, the $|C_{6}|$ coefficients for $|nD_{3/2},m_{j}=3/2\rangle$ of the 87Rb atom with $n=53$, 62, and 82 are measured directly between two isolated single Rydberg atoms separated by a controlled distance of a few micrometers,Béguin et al. (2013) and the angular dependence of the effective interaction energy of the two $82D_{3/2}$ ($82S_{1/2}$) atoms is also demonstrated.Barredo et al. (2014) Recently, with the advancement of experimental techniques enabling the precise arrangement of a large number of atoms in defect-free programmable geometric configurations across various dimensions, the vdw-type Rydberg interaction has found applications in simulating large-scale Ising-type quantum spin models,Bernien et al. (2017); Lienhard et al. (2018); de Léséleuc et al. (2018a); Keesling et al. (2019); Ebadi et al. (2021); Scholl et al. (2021); Bharti et al. (2023); Kim et al. (2024) generating and manipulating Schrödinger cat states,Omran et al. (2019) designing frustrated magnets,Glaetzle et al. (2014, 2015) and exploring nonequilibrium physics.Ding et al. (2020); Bluvstein et al. (2021); Su et al. (2023); Shao et al. (2023); Zhao et al. (2024)
II.3 Emergence of Rydberg superatoms
In the fields of chemistry and atomic physics, the concept of a superatom is used to characterize a cluster of atoms that exhibits collective behavior and possesses characteristics distinct from its constituents.Knight et al. (1984); Inoshita, Ohnishi, and Oshiyama (1986); Khanna and Jena (1995) These superatoms possess unique electronic structures and reactivity,Luo and Castleman (2014); Reber and Khanna (2017) making them highly compelling for a wide range of applications in materials science and nanotechnology.Castleman Jr and Khanna (2009); Doud et al. (2020); Yu et al. (2023) Similarly, a Rydberg superatom denotes a collectively excited state of an ensemble of $N$ atoms, where the Rydberg blockade inhibits the simultaneous excitation of multiple Rydberg atoms in the ensemble.Jaksch et al. (2000) In a two-body system, this blockade effect can be achieved using either the Förster resonance or vdw interaction, as confirmed by experimental observations. Vogt et al. (2006); Urban et al. (2009); Gaëtan et al. (2009); Zeng et al. (2017) However, in a many-body system, the construction of a Rydberg superatom is typically facilitated by the Rydberg interaction in the form of a pairwise additive vdw potential.Heidemann et al. (2007, 2008); Dudin et al. (2012); Zeiher et al. (2015); Stiesdal et al. (2020) This is particularly true for the Rydberg-$S$ state, which possesses only one repulsive branch in its molecular potential, preventing a breakdown of the Rydberg blockade.Walker and Saffman (2005); Pohl and Berman (2009); Vermersch, Glaetzle, and Zoller (2015); Derevianko et al. (2015); de Léséleuc et al. (2018a) Here, we will provide a concise theoretical analysis of the mechanism and associated properties of Rydberg superatoms. Detailed explanations of the experimental procedures and their applications in quantum information and quantum optics will be discussed in the following sections.
Figure 2: (a) Rydberg blockade mechanism. The double excitation of the Rydberg state $|rr\rangle$ is inhibited within the blockade radius $R_{b}$ owing to the large energy shift induced by the strong repulsive interaction. (b) Schematic view of a Rydberg superatom comprised of $N$ atoms sharing a single Rydberg excitation enclosed within a blockade sphere.
Consider an ensemble of $N$ two-level atoms, with ground state $|g\rangle$ and Rydberg state $|r\rangle$, being illuminated by a resonant laser field characterized by a Rabi frequency $\Omega_{0}$. In the interaction picture, the Hamiltonian of the system reads
$\hat{H}_{I}=\sum_{i=1}^{N}\frac{\hbar\Omega_{0}}{2}|g_{i}\rangle\langle r_{i}|+{\rm H.c.}-\frac{1}{2}\sum_{i\neq j}\frac{C_{6}}{R_{ij}^{6}}|r_{i}r_{j}\rangle\langle r_{i}r_{j}|,$ (15)
where a negative value of $C_{6}$ denotes repulsive vdw interactions, while a positive value signifies attractive interactions. At the same time, for the sake of simplification, we temporarily disregard the impact of atomic position and motion on the relative phase between individuals. Under the influence of Eq. (15), if all atoms in the ensemble are initially set in the ground state $|G\rangle=|g_{1},g_{2},...,g_{N}\rangle$, they will collectively transition to a symmetric single atom excitation state, assuming $C_{6}/R_{ij}^{6}\gg\hbar\Omega_{0}$. When examining the simple case of the collective excitation of two atoms, as illustrated in Fig. 2(a), the associated collective blockade radius is defined as $R_{b}=[C_{6}/(\hbar\sqrt{2}\Omega_{0})]^{1/6}$,Honer et al. (2011) where the interaction energy equals the linewidth of excitation dominated by the power broadening $\sqrt{2}\Omega_{0}$. Within this range, Rydberg atoms manifest a robust repulsive interaction, constraining excitation to a single excited symmetric state $|\psi_{+}\rangle=(|gr\rangle+|rg\rangle)/\sqrt{2}$ for the pair of atoms, assuming both atoms are driven with the same Rabi frequency. An offset in Rabi frequencies of two atoms may freeze the dynamics in a single atom, inhibiting the creation of the collective single excited state.Srivastava, Niranjan, and Nath (2019) Expanding the system to the collective excitation of $N$ atoms, the collective blockade radius should be calculated as $R_{b}=[C_{6}/(\hbar\sqrt{N}\Omega_{0})]^{1/6}$, and the corresponding collective state is represented by
$|R\rangle=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle,$ (16)
where all atoms share a single Rydberg excitation. This form of collective excitation also exists in the single-photon superradiance model, where identical ground state atoms interact with a single photon.Scully et al. (2006); Svidzinsky, Chang, and Scully (2008); Scully (2009, 2015); Wang et al. (2015) Figure 2(b) presents a schematic representation of a Rydberg superatom, which is collectively coupled to the ground state $|G\rangle$ with a coupling strength of $\sqrt{N}\Omega_{0}$. We note that the Rydberg superatom is itself a multiparticle $W$ state with maximum entanglement.Dür, Vidal, and Cirac (2000) This entangled state holds significant relevance in quantum information processing.Cabello (2002); Horodecki et al. (2009); Kim et al. (2022) Recent experiments has demonstrated its robustness to both particle loss and external electric field noise, particularly as a logical qubit.Xu et al. (2021); Spong et al. (2021) Therefore, the emergent properties of a Rydberg superatom are derived from the collective behavior of its constituent atoms rather than from the properties of individual atoms alone. These properties encompass enhanced nonlinear optical responses and strong light-matter coupling, both of which hold promise for applications in quantum information processing and quantum optics. As the Rydberg superatom system progresses to the forefront of quantum science, it demonstrates the potential to revolutionize various technological domains, including quantum computing and advanced materials engineering. One important theme in the ongoing research is to develop experimental techniques and explore theoretical models to understand and manipulate properties and dynamics of Rydberg superatoms.
In the context of the Rydberg atom system, along with the Rydberg blockade, phenomena such as the Rydberg antiblockade,Ates et al. (2007); Amthor et al. (2010); Bai et al. (2020a) Rydberg dressing,Henkel, Nath, and Pohl (2010); Pupillo et al. (2010); Honer et al. (2010); Johnson and Rolston (2010); Li, Hamadeh, and Lesanovsky (2012); Balewski et al. (2014); Jau et al. (2016) ground state blockade,Shao et al. (2017); Cidrim et al. (2020) and unconventional Rydberg pumping are encountered.Li and Shao (2018); Shao (2020) While these approaches diversify how Rydberg atoms interact, they fall outside the purview of Rydberg superatoms and, therefore, lie beyond the scope of our discussion here.
III EXPERIMENTAL TECHNIQUES AND ADVANCES
III.1 Coherent laser excitation of Rydberg atoms
An essential step in Rydberg experiments is the preparation of Rydberg atoms. The utilization of lasers offers high-resolution, coherent, and efficient excitation of Rydberg states. After the invention of tunable dye lasers, Rydberg laser spectroscopy with resolution $\sim 0.1\rm~{}nm$ was demonstrated in 1973.Bradley et al. (1973); Stebbings and Dunning (1983) The spectral resolution was improved to $\sim$ 10 GHz in 1979.Zimmerman et al. (1979) In 2003, Rydberg Autler-Townes spectra with MHz resolution were studied with narrow linewidth lasers. Teo et al. (2003)
Figure 3: Different Rydberg laser excitation schemes for 87Rb atom. 297 nm UV laser can couple the ground state $5S$ and the Rydberg state $nP$. 780 nm (420 nm) and 480 nm (1012 nm) lasers can excite atoms from $5S$ to $nS$ or $nD$ via an intermediate state $5P$ $(6P)$ with intermediate detuning $\Delta_{e}$. Due to the finite lifetime of the intermediate states, atoms in these states decay with a rate of $\Gamma_{e}$, which is one of the main sources of error in the two-photon excitation scheme. Using the intermediate states $5P$ and $6S$, atoms can be excited to the $nP$ Rydberg states using the 780 nm, 1367 nm, and 743 nm lasers.
Figure 4: Laser setup of 480 nm and 780 nm lasers used in 87Rb Rydberg excitation. The 780 nm and 960 nm lasers are obtained from external cavity diode lasers (ECDLs). Both lasers are frequency locked to a high finesse, ultra-stable cavity using the Pound-Drever-Hall (PDH) locking technique, which generates “servo bump” due to the limited stabilization bandwidth. Frequency doubling of the power-amplified (using the tapered amplifier) 960 nm laser in the second harmonic generator (SHG) generates blue light at 480 nm. The 780 nm and 480 nm laser beams excite the atoms in a counterpropagating configuration to minimize the Doppler effect.
The laser fields used for Rydberg excitation must be intense enough to guarantee a fast excitation within the Rydberg lifetime and coherence time. The natural lifetimes of high-lying Rydberg states are mainly limited by spontaneous emission and black-body radiation, and can last more than hundreds of microseconds. However, the ground-Rydberg coherence times are usually about a few to tens of microseconds, limited by many decoherence channels, such as the Doppler effect, density-dependent phase shift, fluctuating background electric fields, etc. Therefore, the Rabi frequencies of the excitation laser have to be large, usually on the orders of MHz, to ensure coherent excitations. Such large Rabi frequencies require intense light sources for highly excited Rydberg states since the dipole matrix elements between the ground and Rydberg states decrease with the principal quantum number as $n^{*-1.5}$. In Rydberg tweezer array experiments, it has been shown that a larger Rabi frequency is crucial for high-fidelity operations within the finite ground-Rydberg coherence.Madjarov et al. (2020) In a Rydberg superatom containing $N$ atoms, the effective ground-Rydberg Rabi frequency is improved by a factor of $\sqrt{N}$ due to collective enhancement effects.
Given the considerable energy gaps between the ground and Rydberg states, direct ground-to-Rydberg transition frequencies typically fall within the ultra-violet (UV) range.Wang et al. (2017) Taking Rubidium atoms as an example, lasers with $\sim$ 297 nm wavelength can excite atoms from the $5S$ ground state to the $nP$ Rydberg states.Thoumany et al. (2009) Coherent manipulation between the ground and Rydberg states using a UV laser was demonstrated in Fig. 3, see Ref. Hankin et al., 2014. The single photon UV scheme is ideal for realizing the Rydberg dressing and has also been used in the ground-Rydberg state quantum memory.Li and Kuzmich (2016)
The choice of the excitation scheme depends on the experimental aim. A commonly adopted scheme involves an intermediate state and uses two-photon Rydberg excitation, which can benefit from electromagnetically induced transparency (EIT). With the 780 nm-480 nm (420 nm-1012 nm) two-photon scheme, the ground state Rb atom can be excited to the $nS$ or $nD$ Rydberg states, as shown in Fig. 3. Figure 4 shows the laser setup for the 780 nm-480 nm scheme. The 780 nm light field can be easily obtained from a commercially available external cavity diode laser (ECDL), while the 480 nm light comes from a second-harmonic-generator (SHG) seeded by power-amplified 960 nm laser light. To achieve a strong blockade effect, the laser linewidth must be narrow enough compared to the Rydberg interactions. Therefore, both 780 nm and 960 nm ECDLs are frequency stabilized to an ultralow expansion cavity with high finesse to guarantee narrow linewidth. However, the Pound-Drever-Hall (PDH) laser locking technique leads to large “servo bumps” outside the stabilization bandwidth, which could cause severe decoherence during laser excitation.
[…]
The Rydberg excitation fidelity is affected by various decoherence and loss mechanisms. For the single-photon scheme, a major source of errors in coherent excitation between ground and Rydberg states is the Doppler effect due to atomic motions, which results in a random detuning $\delta=kv$, where $v$ is the velocity of the atom along the wave vector of the laser beam $k$. Limited by the available UV laser power and the small dipole moment between the ground and Rydberg states, the Rabi frequency of the single-photon excitation $\Omega_{\rm UV}$ is on the order of a few MHz. As shown in Fig. 5(a), the ground-Rydberg states $\pi$-pulse excitation error exceeds a few percent for an atomic temperature of ${50}~{}{\mu}$K and $\Omega_{\rm UV}\sim$ MHz. Infidelity caused by the Doppler effect can be suppressed by lowering the temperature or increasing the Rabi frequency.
Compared to the UV laser in single-photon excitation, it is easier to achieve high power for the visible and near-infrared lasers used in two-photon schemes. Therefore, the Rabi frequency of the ground-Rydberg state $\Omega=\Omega_{1}\Omega_{2}/(2\Delta_{e})$ in two-photon excitation can reach tens or even hundreds of MHz, where $\Omega_{1}$ $(\Omega_{2})$ is the Rabi frequency for the ground-intermediate (intermediate-Rydberg) transition, and $\Delta_{e}$ is the intermediate-state detuning, as shown in Fig. 3.
In the two-photon scheme, if a counterpropagating excitation protocol is employed, the Doppler effect will be less severe than in the single-photon case. As shown in Fig. 4, by using counterpropagating excitation beams, the Doppler effect can be minimized.Levine et al. (2018) It is also obvious that the 780 nm-480 nm excitation scheme has a smaller Doppler-induced error than that of the 420 nm-1012 nm scheme. However, regardless of the excitation approach used, achieving a Doppler-induced error well below 1% is achievable by either lowering the atomic temperature or applying a strong laser (large Rabi frequency).
[…]
Figure 5: Influence of Rabi frequency ($\Omega_{1}=\Omega_{2}$ for two-photon excitation schemes) on Doppler-induced Rydberg excitation error in different schemes for an atomic temperature of (a) ${50}~{}{\mu}$K and (b) ${5.2}~{}{\mu}$K. The simulation is performed with intermediate detuning $\Delta_{e}$ = $2\pi\times 5687$ MHz in 780 nm-480 nm scheme and $2\pi\times 7.6$ MHz in 420 nm-1012 nm scheme.
Figure 6: Filtering Cavity and feedforward techniques for suppressing the “servo bump”. (a) The frequency-stabilized laser is sent to a narrow bandwidth optical cavity. The “servo bump” phase noise in the transmitted laser will be strongly suppressed, which is used for the injection lock of another laser diode (LD). (b) The phase information of the laser is obtained by a phase detector, which is then feedforward to an electro-optic modulator (EOM) to cancel the phase noise of the laser. The delay fiber is used to minimize the inherent delay time $\tau$ of the phase detector.
As illustrated in Fig. 5, at $50~{}{\mu}$K and $5.2~{}{\mu}$K, we conduct a comparative analysis of the errors using three different laser excitation schemes with varying Rabi frequencies. For the 780 nm-480 nm excitation scheme, the intermediate state detuning is set at $2\pi\times 5687$ MHz,Fu et al. (2022) while for the 420 nm-1012 nm, it is $2\pi\times 7.6$ GHz. Evered et al. (2023) From the simulation, it is evident that Doppler-induced infidelity can be effectively suppressed by increasing the excitation Rabi frequency.
In addition, laser phase noise and intensity noise also contribute to the Rydberg excitation error. Depending on the laser phase stability and the Rabi frequency, the contribution of phase noise to the excitation error could be more than a few percent de Léséleuc et al. (2018b) without suppressing the servo bump. By using an optical cavity as a filter, as shown in Fig. 6(a), the servo bump, which is typically around MHz, is strongly suppressed by the cavity since its transmission bandwidth can be narrower than kHz. Levine et al. (2018) Moreover, Fig. 6(b) shows an alternative protocol, the feedforward technique, for suppressing the servo bump. A phase discrimination circuit based on the beat signal of the laser Li et al. (2022a) and the residual PDH error signal is used as the phase detector. Chao et al. (2023) The signal obtained from the phase detector is then fed forward to an electro-optical modulator (EOM) with optimized modulation depth so the phase noise of the laser can be reduced. An appropriate length of delay fiber is used to overcome the frequency deviations caused by the inherent delay time of the phase detector. It is reported that after suppressing the servo bump, the fidelity after $9\pi$ pulse  Li et al. (2022a) and the coherence time Levine et al. (2018) can be significantly increased. The intensity noise largely depends on the lasers themselves. For example, the 297 nm laser used in the single-photon scheme in 87Rb atoms can be obtained from the fourth harmonic generation seeded by a power-amplified 1188 nm laser.Li et al. (2019b) Using stimulated Raman scattering fiber amplifier technology, the amplified 1188 nm laser usually has a peak-to-peak intensity noise of a few percent. In addition, the cavities for frequency quadrupling will introduce additional intensity noise in the cavity locking. Considering the above noise sources, the resulting Rydberg excitation errors could be as large as a few percent.
III.2 Preparation of Rydberg superatoms: Many-body Rabi oscillation
The concept of the Rydberg superatom can be connected to the original idea proposed by Dicke, where $N$ atoms are coupled to a single mode of the electromagnetic field at small interatomic distances. In the Dicke model, the collective behavior of closely spaced atoms leads to phenomenon such as superradiance, where the atoms act coherently as a single quantum system.Dicke (1954) This concept can be extended to the Rydberg superatom, where the interatomic distances can be much larger than the wavelength of the excitation light. Despite the larger distances, the atoms in a Rydberg superatom can still form collective excitation states, allowing them to behave as a single, coherent entity. Such a collective effect could lead to a many-body Rabi oscillation with an enhanced Rabi frequency of $\sqrt{N}$$\Omega$.Rabi et al. (1939); Cummings and Dorri (1983); Kumlin et al. (2023) In 2001, Lukin et al. suggested that if an atomic ensemble is confined within the blockade radius,Lukin et al. (2001) many-body Rabi oscillation could be achieved by performing laser excitation between the ground state |$\ket{G}$ and the Rydberg superatom state |$\ket{R}$ of Eq. (16). Coherent many-body Rabi oscillation is not only conclusive evidence for the successful preparation of the Rydberg superatom but also the cornerstone for many Rydberg-atom-related quantum applications, such as deterministic quantum photonic state preparation and manipulation, quantum networks, and quantum simulations.
Figure 7: (a) Rydberg blockade between two single atoms. Reproduced with permission from Urban et al., Nature Physics 5(2), 110-114 (2009). Copyright 2009 Springer Nature.Urban et al. (2009) (b) Rabi oscillations enhanced by two atoms. Reproduced with permission from Gaëtan et al., Nature Physics 5(2), 115-118 (2009). Copyright 2009 Springer Nature.Gaëtan et al. (2009)
Although Rydberg interactions were already observed in the 1980s, Rydberg excitation blockade between two single atoms was not demonstrated until 2009.Urban et al. (2009); Gaëtan et al. (2009) As illustrated in Fig. 7(a), the unfilled circles show the ground-Rydberg state Rabi oscillation of a single atom confined in $\mu$m sized far-off-resonance trap. When a control atom, which is $\sim$ 10 $\mu\rm m$ apart, is driven to the Rydberg state, the excitation of the target atom (filled circles) is suppressed due to a strong two-atom interaction shift of about 9.5 MHz. Urban et al. (2009) This conditioned Rydberg excitation blockade mechanism paves the way for the two-qubit quantum logic gate with neutral atoms.Isenhower et al. (2010)
Figure 7(b) demonstrates the first two-atom enhanced ground-Rydberg Rabi oscillation.Gaëtan et al. (2009) To achieve a strong Rydberg blockade, two atoms are brought within a distance of 4 $\mu\rm m$. Careful selection of Rydberg levels results in the Förster resonance, leading to a Rydberg-Rydberg interaction strength $\sim$ 50 MHz. The Rabi oscillation of two atoms (blue squares) is accelerated compared to the case of a single atom (red circles), by a factor of 1.38(3), close to the expected value of $\sqrt{2}$ in the two-atom blockade model.
Figure 8: Many-body Rabi oscillations in atomic ensembles. (a) Coherent many-body Rabi oscillations of a mesoscopic atomic ensemble. Reproduced with permission from Dudin et al., Nature Physics 8(11), 790-794 (2012). Copyright 2012 Springer Nature.Dudin et al. (2012) (b) Rabi oscillation between the ground state and superatom state for various atom number distributions. Reproduced with permission from Ebert et al., Phys. Rev. Lett. 112(4), 043602 (2014). Copyright 2014 American Physical Society.Ebert et al. (2014) (c) Time evolution of the photon signal and the Rydberg population. Reproduced with permission from Paris-Mandoki et al., Phys. Rev. X 7(4), 041010 (2017). Copyright 2017 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Paris-Mandoki et al. (2017)
In 2012, a coherent many-body Rabi oscillation was demonstrated in a mesoscopic ensemble containing a few hundred atoms.Dudin et al. (2012) To achieve blockade over the entire ensemble, the atoms are confined within a region $\sim$ 6 $\mu\rm m$ and a high-lying Rydberg state |$\ket{102S_{1/2}}$ is used. As shown in Fig. 8(a), pronounced many-body Rabi oscillations are observed between the ground state |$\ket{G}$ and the Rydberg superatom state |$\ket{R}$ under different effective atom numbers $N_{e}$. From the oscillation dynamics, the collective enhancement factor can be extracted and compared with the atom number $N_{a}$. The observed $\sqrt{N_{e}}$ enhancement behavior agrees with Dicke’s prediction. The detection of the Rydberg state is accomplished by converting the superatom state into a single photon. Therefore, this work provides the basis for high-quality Rydberg single-photon source and related quantum optics research.
Several experiments conducted on atomic ensembles have also confirmed a similar many-body Rabi oscillation phenomenon. Figure 8(b) shows the many-body Rabi oscillations where the number of atoms ranges from 1 to 15.Ebert et al. (2014) After the preparation of the Rydberg superatom |$\ket{R}$, it can be transferred to a ground state superatom |$\ket{G^{\prime}}$. The above protocol can be repeated to deterministically prepare the atomic Fock state |$\ket{G^{\prime}}^{n}$. Figure 8(c) displays the many-body Rabi oscillation under strong dissipation,Paris-Mandoki et al. (2017) which facilitates the realization of a Rydberg-based single-photon absorber.Tresp et al. (2016)
Figure 9: Many-body Rabi oscillations in atom array experiments. (a) Collective Rabi oscillations in a quantum gas microscope. Reproduced with permission from Zeiher et al., Phys. Rev. X 5(3), 031015 (2015). Copyright 2015 American Physical Society, licensed under a Creative Commons Attribution 3.0 License.Zeiher et al. (2015) (b) Collective Rabi oscillation in two-dimensional atomic arrays. Reproduced with permission from Labuhn et al., Nature 534(7609), 667-670 (2016). Copyright 2012 Springer Nature.Labuhn et al. (2016) (c) Collective Rabi oscillation in one-dimensional atomic arrays. Reproduced with permission from Bernien et al., Nature 551(7682), 579-584 (2017). Copyright 2012 Springer Nature.Bernien et al. (2017)
In addition to atomic ensembles, many-body Rabi oscillations have also been observed in atom array settings. In 2015, Zeiher et al. demonstrated the preparation of a scalable Rydberg superatom in a quantum gas microscope [Fig. 9(a)].Zeiher et al. (2015) The $\sqrt{N}$ scaling in many-body Rabi oscillation was observed from one to a few hundred atoms. Related advances have also been made in reconfigurable tweezer arrays. Figures 9(b) and 9(c) show the many-body Rabi oscillation with high contrast in one- and two-dimensional atomic arrays.Labuhn et al. (2016); Bernien et al. (2017)
III.3 Quantum interface and coherence property
Figure 10: Efficient matter-light quantum interface using Rydberg superatom. (a) A single atom adsorbs and re-emits a single photon in free space. (b) Storage and retrieval of a single photon using Rydberg superatom. The inset shows the relevant atomic levels.
[…]
$\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}\right)\cdot\mathbf{x}_{j}}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle\otimes|0\rangle.$ (17)
Here, $\mathbf{x}_{j}$ is the position of the $j$-th atom, and $\mathbf{k}_{in},\mathbf{k}_{c}$ are the wave vectors of the input photon and the control laser, respectively. $|0\rangle$ is the vacuum state of the photon field. This Rydberg EIT light storage process efficiently converts a photon into a single Rydberg excitation collectively shared among the atoms inside the Rydberg blockade radius.
Within the coherence time of the Rydberg superatom, the stored excitation can be transferred back to a single photon, with a read-out laser field. The read-out laser transfers the collective excitation from the Rydberg superatom state to the intermediate state,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}\right)\cdot\mathbf{x}_{j}}|g_{1},g_{2},...,e_{j},...,g_{N}\rangle\otimes|0\rangle,$ (18)
where $\mathbf{k}_{r}$ is the wave vector of the readout laser. Due to the finite lifetime in the intermediate state $\left|e\right\rangle$, the collective state soon radiates into a new state,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}-\mathbf{k}_{out}\right)\cdot\mathbf{x}_{j}}|g_{1},g_{2},...,g_{j},...,g_{N}\rangle\otimes|1\rangle_{\mathbf{k}_{out}},$ (19)
where the atoms are back in the ground state $\left|g\right\rangle$ and a single photon is emitted into the direction defined by its wave vector $\mathbf{k}_{out}$. The probability of emitting a single photon into a certain mode $\mathbf{k}_{out}$ is given bySaffman and Walker (2002)
$P\left(\mathbf{k}_{out}\right)\propto\frac{1}{N}\left|\sum_{j}e^{-i\left(\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}-\mathbf{k}_{out}\right)\cdot\mathbf{x}_{j}}\right|^{2}.$ (20)
Assisted by the collective enhancement of the Rydberg superatom, this emission pattern is highly directional: for mode $\mathbf{k}_{out}=\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}$, each atom shares the same phase, and the emission amplitude is maximally enhanced. In single photon storage-and-retrieval experiments, the control laser and the read-out laser are often chosen to have the same mode $\mathbf{k}_{c}=\mathbf{k}_{r}$, so the retrieved single photon will propagate along the same mode as the input, $\mathbf{k}_{out}=\mathbf{k}_{in}$.
The efficient matter-light interface makes the Rydberg superatom a promising candidate for photonic quantum memory. Unlike their ground state counterpart, Rydberg quantum memories interact strongly, providing new possibilities for realizing deterministic quantum photonic devices, such as conditional phase shifter,Firstenberg et al. (2013); Tiarks et al. (2016); Thompson et al. (2017) entangling gate,Tiarks et al. (2019); Stolz et al. (2022) and entanglement filter.Ye et al. (2023)
Rydberg superatom also differs from a single atom in terms of its coherent properties. As a collective excitation, Rydberg superatom is the superposition of each microstates: $\sum_{j=1}^{N}e^{i\phi_{j}}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle$. Therefore, the coherence of the Rydberg superatom largely depends on how well the relative phases $\phi_{j}$ between each microstates are maintained. Any effect that smears the relative phases will cause superatom decoherence.
[…]
To suppress motional dephasing, an effective technique is to confine the atoms within spatial regions that are much smaller than the spin wave period. For ground states superatom, a one-dimensional optical lattice along the direction of the spin wave can be used to overcome the motional dephasing and extend the coherence time to minutes scale.Dudin, Li, and Kuzmich (2013) However, while red-detuned far-off-resonance optical traps are attractive for ground atoms, they are generally repulsive for Rydberg atoms.Saffman and Walker (2005); Adams, Pritchard, and Shaffer (2020) Therefore, traps that could simultaneously confine ground and Rydberg states are essential for extending motional decoherence times. To this end, optical traps with “magic” wavelengths are proposed Zhang, Robicheaux, and Saffman (2011); Saffman and Walker (2005) and demonstrated. Li, Dudin, and Kuzmich (2013) In ground-Rydberg state insensitive lattices, the Rydberg superatom storage times and many-body Rabi oscillation coherence times have been extended to tens of microseconds. Lampen et al. (2018); Mei et al. (2022)
In addition to the ground-Rydberg-state magic trapping lattice, progress in extending the superatom coherence times was also made using other approaches. Using 297 nm single-photon Rydberg excitation, the quantum state shelving between Rydberg and ground states superatom was demonstrated. Li and Kuzmich (2016) The relatively short-lived Rydberg superatom can be coherently transferred to the ground state superatom, whose coherence time can be orders of magnitude longer. Furthermore, atomic samples prepared at lower temperatures also help to mitigate motional dephasing. Schmidt-Eberle et al. (2020) In addition, it was recently shown that Rydberg spin wave motional dephasing can also be suppressed by introducing another spin wave using off-resonant ac-Stark lattice modulation.Kurzyna et al. (2024)
Due to their collective nature, the coherence properties between two Rydberg superatoms also differ from those of two single atoms. Multiple collective Rydberg excitations (multiple Rydberg superatoms) can be prepared for ensembles separated by a distance larger than the blockade radius. After the preparation of two Rydberg superatoms, the interaction between the pair of Rydberg atoms leads to the accumulation of phases over time,
$|RR\rangle\sim\sum_{j,j^{\prime}\neq j}^{N}e^{i\frac{V_{jj^{\prime}}}{\hbar}t}|g_{1},...,r_{j},...,r_{j^{\prime}},...,g_{N}\rangle.$ (21)
Here, $V_{jj^{\prime}}=C_{6}/x_{jj^{\prime}}^{6}$ is the vdW interaction between the Rydberg atom pair separated by $x_{jj^{\prime}}$. Since the interaction is distance-dependent, the accumulation of random phases over time leads to a collective two-body dephasing.Bariani et al. (2012) To reduce such a dephasing mechanism, the interaction disorder between Rydberg superatoms needs to be suppressed, for instance, by making the superatom size much smaller than the separation between superatoms.
Notably, the Rydberg interaction-induced decoherence mechanism can be harnessed for the effective preparation and manipulation of photonic quantum states.Bariani et al. (2012); Stanojevic et al. (2012) For example, interaction dephasing-based Rydberg single-photon source, Li et al. (2022b); Berman and Kuzmich (2024) non-local quantum optics,Busche et al. (2017) and dissipative entanglement creationYe et al. (2023) have been experimentally demonstrated.
IV APPLICATIONS IN QUANTUM INFORMATION PROCESSING
IV.1 Quantum logic gates
Figure 11: (a) The displayed outcome presents both the results of Monte Carlo simulations and analytical approximations for a random gas in a finite volume, illustrating the probability distribution for frequency shifts of doubly excited Rydberg states. (b) Dipole blockade facilitates the generation of successive Fock states. Reproduced with permission from Lukin et al., Phys. Rev. Lett. 87(3), 037901 (2001). Copyright 2001 American Physical Society.Lukin et al. (2001)
The strong and long-range interactions furnish a direct pathway for the realization of two-qubit gates within Rydberg superatoms. The pioneering gate protocol in Ref. Lukin et al., 2001, entails a Rydberg superatom blockade gate. In an ensemble comprising $N$ identical atoms, each atom has the opportunity to be excited with equal probability, leading to the excitation of symmetric collective states exclusively starting from the ground state. However, the presence of resonant dipole-dipole interactions introduces a spectrum of doubly excited states, each state being separated by an energy gap of approximately $\hbar\bar{\kappa}=\wp_{rp^{\prime}}\wp_{rp^{\prime\prime}}/V$ in ensembles confined within a finite volume $V$, where $r$, $p^{\prime}$, and $p^{\prime\prime}$ denote long-lived Rydberg states, $\wp_{rp^{\prime}}$ and $\wp_{rp^{\prime\prime}}$ denote the dipole matrix elements for the corresponding transitions. Should the energy gap exceed the linewidth of the Rydberg state, resonant excitations from singly to doubly excited states experience significant suppression. Figure 11 illustrates the probability distribution for frequency shifts of doubly excited Rydberg states by Monte Carlo simulations of $3\times 10^{4}$ positions within a rectangular box along with the corresponding analytical approximations for a random gas. The strength of the interaction $\kappa$ between any pair of atoms exceeds $\bar{\kappa}$, dominating most of the probability. This inhibition facilitates the production of collective spin states, enhancing spectroscopic sensitivity beyond the standard quantum limit. It also aids in the preparation of a qubit that exhibits superpositions between the ground state and the single-excited storage state. This capability proves valuable for generating entanglement between two groups of atoms, drawing on principles similar to those described in Ref. Jaksch et al., 2000. This regime sets the groundwork for the emergence of superatoms in atomic gases, encompassing the Rydberg dressing, and opens avenues to scalable quantum logic gates within optical or magnetic trap arrays. Honer et al. (2010); Gaul et al. (2016); Jo et al. (2020) Subsequently, experimental demonstrations of superatom gates have been achieved, Reetz-Lamour et al. (2008) along with the proposal of a scheme combining EIT and Rydberg blockade in an atomic ensemble. Weatherill et al. (2008)
To streamline operational steps and address practical applications, a single-step multiatom entanglement scheme based on individual atoms within an ensemble has emerged. Müller et al. proposed a scheme to achieve a high-fidelity gate of many particles using Rydberg interactions and combining EIT, Müller et al. (2009) which can be applied to multiparticle interferometry measurements. Micheli et al. (2004); Seaman et al. (2007). The basic idea is that a control atom conditionally affects the dynamics of spatially separated ensemble atoms through long-range Rydberg interactions. The control atom consists of two ground states $|0_{c}\rangle$ and $|1_{c}\rangle$, and a Rydberg state $|r_{c}\rangle$, where states $|1_{c}\rangle$ and $|r_{c}\rangle$ are resonantly coupled. Ground states $|A_{e}\rangle$ and $|B_{e}\rangle$ of atoms in the ensemble are off-resonantly coupled to intermediate state $|P_{e}\rangle$ (Rabi frequency $\Omega_{p}$), and to Rydberg state $|R_{e}\rangle$ (Rabi frequency $\Omega_{c}$). Under the EIT condition (two-photon resonance and $\Omega_{c}\gg\Omega_{p}$), the collective $\pi$-pulse evolution can be performed (e.g. $|0_{c}\rangle|A^{N}_{e}\rangle\to|0_{c}\rangle|A^{N}_{e}\rangle$) if the control atom is in state $|0_{c}\rangle$. However, when the control atom is initialized in state $|1_{c}\rangle$ and excited to $|r_{c}\rangle$, causing an energy shift of the $|R\rangle$ state. Using the resulted Raman transfer between $|A\rangle$, and $|B\rangle$, the collective operations $|1_{c}\rangle|A_{e}^{N}\rangle$leftrightarrow $|1_{c}\rangle|B_{e}^{N}\rangle$, and $|1_{c}\rangle|B_{e}^{N}\rangle$leftrightarrow $|1_{c}\rangle|A_{e}^{N}\rangle$ can be performed. Based on this model, the fidelity of the controlled-${\rm NOT}^{N}$ logic gate can reach 99%. Similar setting can also be used for nonadiabatic geometric quantum computation to reduce evolution time and mitigate decoherence. Xu et al. (2012); Zhao et al. (2018); Maity and Purkait (2022); Guo et al. (2020) In 2018, nonadiabatic geometric quantum computation based on superatoms was proposed, Zhao et al. (2018) and can be extended to the 3-qubit case with two species of superatoms in a two-mode cavity. Maity and Purkait (2022) Guo et al. utilized dynamical invariant and zero systematic-error sensitivity methods to reverse design gate pulses, Guo et al. (2020) optimizing gate fidelity and robustness against control errors. The average fidelity of the two-qubit controlled-NOT (CNOT) gate in their scheme is 0.9971.
For multiatom ensemble qubits, Saffman’s group demonstrates coherence control between two ensemble qubits and achieves Rydberg blockade between them with a fidelity of 0.89(1).Ebert et al. (2015) However, fluctuations in the collective Rabi frequency could greatly affect the ensemble qubits operation fidelity. Thus, a scheme is proposed to simulate quantum process tomography to characterize Rydberg gates based on collective states using optimal stimulated Raman adiabatic passage pulses. Beterov et al. (2016b) Recently, significant progress has been made in achieving entanglement gates through coherent manipulation of deterministically assembled arrays, employing local optical control to achieve entanglement using interatomic interactions, although individual addressing of atoms remains a challenge experimentally. Zeng et al. (2017); Scholl et al. (2023); Evered et al. (2023); Ma et al. (2023) Additionally, due to the excellent transmission characteristics of photons but the lack of strong interactions, entangling gate schemes for photon-atom hybridization have been proposed. Garcia-Escartin and Chamorro-Posada (2012); Paredes-Barato and Adams (2014); Hao et al. (2015); Das et al. (2016); Sárkány, Fortágh, and Petrosyan (2015) Specifically, the combination of a Rydberg blockade and optical cavities effectively enhances the gate robustness, Das et al. (2016) while a two-photon interference through virtual photon exchanges renders the internal state dynamics of two neutral-atom qubits insensitive to the thermal photon state. Sárkány, Fortágh, and Petrosyan (2015)
Figure 12: Level scheme for EIT storage of the control qubit (a) and target qubit (b). (c) The schematic of the experimental setup. Here, $|R\rangle$ and $|L\rangle$ are used to define polarization states of photons. Reproduced with permission from Tiarks et al., Nature Physics 15(2), 124-126 (2019). Copyright 2019 Springer Nature.Tiarks et al. (2019)
Tiarks et al. achieved the first experimental realization of a photon-photon quantum gate using Rydberg interactions, as documented in their study.Tiarks et al. (2019) By harnessing both EIT and Rydberg interactions within an ultracold atomic ensemble, they effectively encoded quantum information carried by photons into specific atomic states, as depicted in Fig. 12(a). This mapping facilitates the control of photon loss through post-selection, thereby mitigating its detrimental impact on quantum information and ensuring resilience against errors stemming from photon loss. Notably, the photon interaction occurs exclusively when they share the same polarization $|L\rangle$ due to the Rydberg component, and their interaction during temporal and spatial overlap results in the accumulation of a conditional phase shift, thus realizing a controlled phase flip gate, as illustrated in Fig. 12(b). By altering the basis, this gate can be converted into a CNOT gate. The fidelity of the CNOT truth table and the entangling-gate fidelity, post-selected upon the detection of control and target photons, are reported to be 70(8)and 63.7(4.5)%, respectively. This technique employs effectively Rydberg interactions and EIT, allowing for strong, long-range interactions between photons while also solving the difficulty of restricted photon interactions. By mapping Rydberg interactions onto photons, this gate approach stores and manipulates quantum information, demonstrating high fidelity and potential advances in quantum communication systems and quantum networking applications.
Figure 13: Schematic diagram of converting a circuit into an atomic arrangement, which consists of two types of atoms, and superatoms are used as impurities. Reproduced with permission from Cesa et al., Phys. Rev. Lett. 131(17), 170601 (2023). Copyright 2023 American Physical Society.Cesa and Pichler (2023)
Recently, the Rydberg superatoms were utilized in a model for quantum computation using dual-species Rydberg atom arrays, Cesa and Pichler (2023) which relies solely on global driving [see Fig. 13], thus eliminating the requirement for local addressing of the qubits. Ceas and Pichler introduce two distinct constructions: One imprints the circuit within the trap positions of the atoms, executed by pulses; the other encodes the algorithm within the global driving sequence, making the atom arrangement circuit independent. Central to this scheme is the concept of information flow, specifically focusing on the ability of qubits to propagate described by $|\Psi(k+1)\rangle={\rm U}_{\mathcal{B}}{\rm U}_{\mathcal{A}}{\rm U}_{\mathcal{B}}|\Psi(k)\rangle$, achieved under the conditional occurrence of the Rydberg blockade. Furthermore, the initialization of the circuits must align with the information flow. For the basic requirements of quantum computation, the construction of single- and two-qubit gates requires the assistance of impurity superatoms [see Figs. 13(c) and 13(d)]. Arbitrary single-qubit gates, including the Hadamard gate, are realized by using superatoms $\mathcal{A}$ within the wires to enhance the Rabi frequency, facilitating collective phase gates $\mathcal{Z}_{\rm tot}=\bigotimes_{Q}{\rm Z}_{Q}$. Additionally, for two-qubit entangling gates, another superatom $\mathcal{B}$ placed between the wires can induce a $\rm CZ$ gate. The study demonstrates that a quadratic increase in atom number is sufficient to eliminate the need for local control, thereby achieving a universal quantum processor. The authors also provide explicit protocols for all stages of arbitrary quantum computations and discuss error suppression strategies tailored to their model. Although the scheme is grounded in dual-species processors adhering to Rydberg blockade constraints, it holds potential for adaptation to alternative setups. This study opens new avenues for universal quantum computation with Rydberg superatoms.
IV.2 Quantum entanglement
Quantum entanglement plays a crucial role in quantum computing and quantum information processing. Early experiments of entangled states have focused on photons and ions. Prevedel et al. (2009); Wieczorek et al. (2009); Häffner et al. (2005) However, these schemes face challenges in scaling to multiqubit entanglement and have a low success probability. Rydberg ensembles offer a promising solution to overcome these limitations and realize scalable entanglement of neutral atoms. Atom vapors serve not only as quantum memory, but also as “good” qubits due to strong Rydberg interactions. The theoretical study shows that high-efficiency cluster states can be created via single-qubit measurements with Rydberg ensembles. Zwierz and Kok (2009) Utilizing a fast entanglement generation scheme with a Lipkin-Meshkov-Glick-type Hamiltonian, a 700-atom cat state has been generated with a 50-fold improvement in the size of the entangled state. Khazali (2018) To further enhance fidelity, Gärttner proposes a scheme to herald the entangled state by detecting an ion, achieving near-unity fidelity. Gärttner (2015) Additionally, addressing decoherence concerns, one can combine Zeno dynamics and shortcuts to adiabaticity with Rydberg superatoms, encoding quantum information in the collective state of superatoms to rapidly realize entangled states, including Bell states and $W$ states. Ji et al. (2020) Similarly, fast conversion of three-particle Dicke states to four-particle Dicke states with Rydberg superatoms can be implemented based on Zeno dynamics and shortcuts to adiabaticity, offering high fidelity and robustness against decoherence. Liu et al. (2023b) In Rydberg ensembles, large-scale entangled states can also be prepared theoretically by sending two Rydberg superatoms belonging to two different atomic $W$ states into a cavity without qubit loss. Zhang et al. (2022a)
Figure 14: Entanglement between Rydberg superatom and photon. (a) Schematic diagram of the entanglement of light with optical atomic excitation. Mapping the collective atomic excitation into quantum fields $|\phi\rangle_{A,B}$ by implementing resonance on the $|r\rangle$ and $|e\rangle$ used laser fields $\Omega_{A,B}$. Reproduced with permission from Li et al., Nature 498(7455), 466-469 (2013). Copyright 2013 Springer Nature.Li, Dudin, and Kuzmich (2013) (b) Experimental preparation of semideterministic atom-photon entanglement. Reproduced with permission from Li et al., Phys. Rev. Lett. 123(14), 140504 (2019). Copyright 2019 American Physical Society.Li et al. (2019c)
Photons are excellent carriers for quantum information, and the entanglement between atom and photon holds promise for achieving long-range quantum information distribution. Most ensemble-based atom-photon entanglement schemes are probabilistic and suffer from limited scalability. In 2013, Li et al. introduced a new protocol to generate entanglement between light and optical atomic excitations with intrinsic determinism. Li, Dudin, and Kuzmich (2013) In this scheme, the dephasing of optical atomic coherence is suppressed by state-insensitive confinement of ground and Rydberg states of ultra-cold atomic gas trapped in a one-dimensional optical lattice at 1004 nm. This confinement eliminates the differential energy shift between the ground state and the Rydberg state $\delta U=U_{r}(r)-U_{g}(r)$ and maintains coherence between them. Such a scheme has significant implications for extending the ground-Rydberg coherence and building functional Rydberg quantum networks. Figure 14(a) illustrates the three-step process for the entanglement protocol: (i) Excitation of the atomic ensembles from the ground state $|G\rangle$ to the single excited Rydberg state $|R\rangle$. (ii) Generation of the entanglement state $|R\rangle\rightarrow|R\rangle|0\rangle_{A}+|G\rangle|1\rangle_{A}$ through a retrieval field $\Omega_{A}$. (iii) Verification of the atom-photon entanglement by analyzing the correlations of photoelectric detection events. This scheme is compatible with quantum logic operations and long-term storage of quantum information, paving the way for quantum networks that enable secure communication and distributed quantum computing.
In 2019, an alternative atom-light entanglement protocol was experimentally demonstrated, leveraging Rydberg blockade and Raman coupling techniques. Li et al. (2019c) The manipulation of atoms within a confined volume of approximately 10 $\mu$m in all three dimensions ensures high effectiveness of Rydberg blockade, with an atomic density of approximately $10^{11}/\rm cm^{3}$. In the schematic depicted in Fig. 14(b), the atomic ensemble is initially prepared in a ground state $|g\rangle$, following which Rydberg blockade is employed to sequentially generate two collective excitations—one in a ground state $|s\rangle$ and the other in a Rydberg state $|r\rangle$. These collective excitation states, referred to as two superatoms, possess distinct momenta, are then entangled by incorporating the momentum degree of freedom and interfering them through Raman coupling. Finally, one excitation is retrieved using a read beam, converting the momentum-entangled atomic state into entanglement between the polarization of the single photon and the momentum of the remaining atomic excitation, thereby achieving atom-photon entanglement. The high fidelity measurement of entanglement between the polarization of the single photon and the momentum of the atomic excitation underscores the practicality of employing this approach for quantum information processing tasks like quantum repeaters and networks. Although this scheme is semi-deterministic compared to the approach in Ref. Li, Dudin, and Kuzmich, 2013, it remains immune to phase fluctuations and is devoid of high-order or vacuum components, which are essential for remote entanglement generation. After that, the same group presented a report detailing the deterministic creation of entanglement between an atomic ensemble and a single photon.Sun et al. (2022) This achievement was accomplished through a cyclical process termed “retrieving and patching”, which was applied sequentially to the Rydberg levels. By leveraging the Rydberg blockade effect, they ensured that a full excitation was generated only if a preceding excitation existed in one of the Rydberg levels. This process led to the creation of a photon in a delayed temporal mode, entangled with the Rydberg levels hosting the collective excitation. Verification of the entanglement between the atomic ensemble and the single photon was conducted by retrieving the atomic excitation as a second photon and performing correlation measurements. The results yielded an entanglement fidelity of 87.8%.
The entanglement between photons holds tremendous potential across various fields, such as metrology and quantum computing. A challenge is that interactions between photons are inherently weak. A lot of efforts have been spent on creating sizable photon-photon interactions. Heisenberg and Euler (1936); Busche et al. (2017); Firstenberg et al. (2013) Numerous schemes have been proposed for the preparation of multiphoton entanglement, including spontaneous parametric downconversion Pan et al. (2012); Zhong et al. (2018) and the entangling of single photons through atom-photon interactions. Gheri et al. (1998); Schön et al. (2005) However, these schemes often suffer from drawbacks such as poor scalability of the photon number, complex setups, and other requirements on the system. Recently, Yang et al. proposed a novel scheme to generate multiphoton entanglement with a Rydberg superatom. Yang et al. (2022a) In their approach, an atomic ensemble is used, where an atomic entanglement state $|\Psi\rangle_{\mathrm{a}}$ is prepared by applying collective pulses with the Rydberg blockade mechanism. Subsequently, a “retrieving and patching” sequence is applied to $|r_{1}\rangle$, causing the Rydberg superatom to emit a photon and recreate the excitation. This sequence is repeated for $|r_{2}\rangle$, resulting in the creation of the entanglement state $(|0\rangle_{1}|1\rangle_{2}|L\rangle+|1\rangle_{1}|0\rangle_{2}|E\rangle)/\sqrt{2}$. By repeating these processes $(m-1)$ times and retrieving the two atomic qubits, the GHZ-type multiphoton entanglement $(|L\rangle^{\otimes m}+|E\rangle^{\otimes m})/\sqrt{2}$ can be obtained, where $|L\rangle$ and $|E\rangle$ represent a photon in the late and early modes, respectively. To enhance the performance of multiphoton entanglement, the scheme employs a low-finesse cavity to improve photon collective efficiency and polarizing beamsplitters with a Pockels cell in measurement to minimize exposure. The fidelity of the final prepared three- and six-photon GHZ states is $82.9\pm 0.3\%$ and $61.8\pm 2.6\%$, respectively.
Strongly interacting Rydberg superatoms play a crucial role not only in generating entanglement but also in enabling deterministic manipulation of entangled photonic states. In Ref. Ye et al., 2023, a novel photonic entanglement filter was realized using two Rydberg superatoms. The demonstrated entanglement filter transmits the desired photonic entangled state and blocks unwanted ones. As illustrated in Fig. 15(a), to achieve quantum state filtering, the undesired photonic components, $|HH\rangle$ and $|VV\rangle$ states, are converted into double Rydberg excitations in the same atomic ensemble via EIT photon storage. These double excitations can be eliminated using either Rydberg blockade or interaction-induced dephasing. On the other hand, the target entangled state, $|H\rangle|V\rangle+|V\rangle|H\rangle$ is protected in a decoherence-free subspace and can pass through the entanglement filter. Compared to prior repeat-until-success approaches based on probabilistic linear-optical systems, the Rydberg entanglement filter harnesses the strong atomic interaction and is, therefore, inherently deterministic. As shown in Fig. 15(b), photonic entanglement with fidelity $\sim 99\%$ can be extracted from an input state even when the initial fidelity is low. Such an entanglement filter opens new possibilities for scalable photonic quantum information processing with Rydberg superatom arrays.
Figure 15: Photonic entanglement filter with Rydberg superatoms. (a) Illustration of Rydberg entanglement filter protocol. (b) Extracting high-fidelity entangled state from input state with low fidelity. Reproduced with permission from Ye et al., Nature Photonics 17(6), 538–543 (2023). Copyright 2023 Springer Nature.Ye et al. (2023)
The overall outlook for creating quantum entanglement using Rydberg superatoms is very positive. Areas including quantum computing, quantum communication, and basic quantum physics stand to benefit from further research and development in this field. These upgrades might pave the way for ground-breaking discoveries with far-reaching consequences in various science and technology areas.
IV.3 Quantum simulation
Quantum simulation serves as a powerful tool for understanding and predicting many-body dynamics, allowing access to a wide range of phenomena, including thermal equilibrium relaxation in closed systems and non-thermal behaviors in integrable and disordered systems.Daley et al. (2022) A particularly promising platform for investigating the dynamics of strongly interacting many-body quantum systems is provided by coherent laser-driven Rydberg atomic gases.Weimer (2015); Lechner and Zoller (2015); Ding et al. (2024) Currently, spin systems based on Rydberg atoms have emerged as a promising platform for quantum simulation because of their superior interaction flexibility and strength, high controllability, and effective isolation from the environment.
One theoretical strategy suggested by Sun and Robicheaux involves simulating Rydberg gases and exploring two-body correlation by driving the superatoms, which can be subdivided into smaller regions called pseudoatoms, with a laser having a Rabi frequency following a Poissonian distribution. Sun and Robicheaux (2008) This method incorporates the classical correlationZhou et al. (2006) and provides a more accurate depiction of the Rydberg gas compared to the mean field approximation. Using the superatom picture, Olmos and colleagues conducted extensive theoretical research on fermionic collective excitations and thermalization in a symmetric quasi-one-dimensional ring lattice gas of Rydberg atoms. Olmos, González-Férez, and Lesanovsky (2009); Olmos, Müller, and Lesanovsky (2010) In the theoretical studies of Refs. Pohl, Demler, and Lukin, 2010; van Bijnen et al., 2011, the Rydberg superatoms serve as building blocks for the creation of Rydberg crystals using tailored laser excitation schemes. These crystals are used as representations of dilute metallic solids with tunable lattice parameters, allowing access to a wide range of fundamental phenomena in condensed matter physics.
Figure 16: Average density of holes $\rho_{h}$, Mandel $Q$ parameter for the total number of holes, and amplitudes of the density–density correlation function $g^{\left(2\right)}\left(d\right)$ for the periods of $d=2a$ (blue circles) and $d=3a$ (red circles), versus the hopping rate $\Gamma_{t}$. Reproduced with permission from Letscher et al., New J. Phys. 19(11), 113014 (2017). Copyright 2017 Institute of Physics, licensed under a Creative Commons Attribution 3.0 licence.Letscher, Petrosyan, and Fleischhauer (2017)
[…]
Figure 17: Ground state population and corresponding Fourier transform in $N$-body ($N=3,4$) configuration. Reproduced with permission from Kim et al., PRX Quantum 1(2), 020323 (2020). Copyright 2020 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Kim et al. (2020)
The scheme outlined in Ref. Kim et al., 2020 suggests the continuous tuning of the quantum Ising Hamiltonian of Rydberg atoms arranged in three dimensions. In this arrangement, the atoms and pairs of Rydberg blockaded atoms are depicted as vertices and edges, respectively, as illustrated by the geometric red configuration in Fig. 17. Consequently, various connected graphs of $N$ body configurations of Rydberg atoms are constructed, and their eigenenergies, along with different geometric intermediates, are explored during the structural transformation process. The ground state population and the corresponding Fourier transform of the configurations in the superatom regime are depicted in Fig. 17 for $N=3,4$. Remarkably, the experimentally measured graph-dependent eigenspectra (represented by filled circles) align well with the model calculations of the few-body quantum Ising Hamiltonian (illustrated by lines). This high-dimensional programming of qubit connectivity demonstrated in the study represents a significant step toward the application of programmable quantum simulators.Byun, Kim, and Ahn (2022); Jeong et al. (2023)
Figure 18: The numerically calculated probabilities $q_{Ry}\left(n\right)$ to excite $n$ Rydberg atoms in the mesoscopic atomic ensemble with $\bar{N}=7$ atoms, randomly distributed in the optical dipole trap with different radius $r$. Reproduced with permission from Beterov et al., Phys. Rev. A 90(4), 043413 (2014). Copyright 2014 American Physical Society.Beterov et al. (2014)
One of the challenges in observing the Rydberg blockade is to obtain a precise measurement of the quantity of Rydberg atoms. Due to low detection efficiencies, accurately determining the number of excited Rydberg atoms becomes problematic, which is essential for exploiting the Rydberg blockade in quantum information applications. Beterov et al. investigated the complex dynamics of Rydberg excitations in atomic ensembles influenced by a classical electromagnetic field.Beterov et al. (2014) They demonstrated the observation of collapses and revivals in the average count of Rydberg excitations, resembling the behavior predicted by the Jaynes-Cummings (JC) model. Through the analysis of randomly loaded optical dipole traps, they predicted the occurrence of collapses and revivals in Rabi oscillations, emphasizing how the oscillation frequency depends on the number of interacting atoms. Moreover, the study examines the impact of finite interaction strengths and laser linewidth on the detectability of the revivals. Considering a randomly loaded optical dipole trap with an average atom number of $\bar{N}=7$, data in Fig. 18 illustrate the excitation dynamics in a Rydberg state in the ensemble. In Fig. 18(a), the simulated dynamics of the probability of single-atom Rydberg excitation $q_{Ry}\left(1\right)$ for $r=2~{}\mu\rm m$ show collapses and revivals, while $q_{Ry}\left(2\right)$ approaches zero, indicating an effective Rydberg blockade. However, with increasing radius of the optical dipole trap, the revivals in $q_{Ry}\left(1\right)$ are diminished, and the probability of exciting two Rydberg atoms $q_{Ry}\left(2\right)$ increases, suggesting the breakdown of the Rydberg blockade. These results provide valuable insights into Rydberg blockade phenomena, reducing the need for precise quantification of Rydberg atoms and offering potential applications in quantum information processing and the investigation of sub-Poissonian atom-number fluctuations in mesoscopic atomic ensembles.
It is noted that significant efforts have been directed towards quantum simulation utilizing Rydberg atom arrays, driven by their exceptional controllability and versatility.Browaeys and Lahaye (2020); Kaufman and Ni (2021) These systems facilitate precise manipulation of individual atoms through laser fields, allowing for the creation of customized potential landscapes and interaction patterns.Omran et al. (2019); Labuhn et al. (2016) Furthermore, the strong and adjustable interactions among Rydberg atoms enable the emulation of various quantum phase transitions.de Léséleuc et al. (2019); Kanungo et al. (2022); Chen et al. (2023a); Bernien et al. (2017); Keesling et al. (2019) The scalability of these arrays, capable of accommodating hundreds of and even moreatoms,Ebadi et al. (2021); Scholl et al. (2021); Bluvstein et al. (2021) offers researchers a robust platform for exploring large quantum systems and addressing computationally challenging problems.Barredo et al. (2016); Endres et al. (2016); Glaetzle et al. (2017); Barredo et al. (2018); Hollerith et al. (2019); Bekenstein et al. (2020); Semeghini et al. (2021); Ebadi et al. (2022); Wu et al. (2022); Bluvstein et al. (2022); Kim et al. (2022); Scholl et al. (2023); Schine et al. (2022); Yan et al. (2022); O’Rourke and Chan (2023); Srakaew et al. (2023); Bornet et al. (2023); Chen et al. (2024) Leveraging unique properties of Rydberg superatoms, we anticipate opportunities for simulating quantum spin models, probing topological phases, and studying thermalization. This can be achieved using techniques such as trapping many atoms per blockade radius in optical lattices or arrays of microscopic dipole traps spaced a few micrometers apart.Côté et al. (2006); Weimer et al. (2010); Viteau et al. (2011); Anderson, Younge, and Raithel (2011); Weber et al. (2015); Gross and Bloch (2017); Whitlock, Glaetzle, and Hannaford (2017); Mei et al. (2022)
V APPLICATIOINS IN QUANTUM OPTICS
In this section, we first present a brief review of recent experimental and theoretical progress focusing on the nonlocal nonlinear responses, as induced by the dipole-dipole interactions and related to the dipole blockade effect, of cold Rydberg atoms driven into the EIT regime.Fleischhauer, Imamoglu, and Marangos (2005); Finkelstein et al. (2023) Then we look back at some important proposals and schemes for achieving efficient nonlinear quantum opticsChang, Vuletić, and Lukin (2014) by considering hybrid systems with Rydberg atoms restricted in cavities, placed close to waveguides, or trapped above atomic chips.
Figure 19: (a) Schematic of a Rydberg EIT experiment with cold 87Rb atoms driven by a probe field on the lower $|g\rangle\leftrightarrow|e\rangle$ transition and a coupling field on the upper $|e\rangle\leftrightarrow|r\rangle$ transition in the ladder configuration. (b) Nonlinear EIT spectra of the probe field where the transmission peak on resonance ($\Delta_{p}=0$) reduces as probe Rabi frequency $\Omega_{p}$ increases. The experiment is done for the $|r\rangle=|60S_{1/2}\rangle$ state at atomic density $\rho=1.2\pm 0.1\times 10^{10}$ cm-3. Reproduced with permission from Pritchard et al., Phys. Rev. Lett. 105(19), 193603 (2010). Copyright 2010 American Physical Society.Pritchard et al. (2010)
Figure 20: (a) A ladder scheme of three-level atoms that interact with a probe and a coupling field. $\mathcal{V}_{\rm vdW}$ denotes the vdW interaction between two atoms in the Rydberg state $|r\rangle$. (b) Truncated level scheme of a superatom composed of $n_{SA}$ atoms, with collectively enhanced transition amplitudes involving different numbers of probe and coupling photons. (c) Top: Probe field transmission against detuning $\Delta_{p}$ for different input probe Rabi frequencies. The thin lines are experimental curves from Ref. Pritchard et al., 2010 while thick lines are numerical simulations based on the superatom model. Bottom: Corresponding intensity correlation function of the transmitted probe field at the exit from the medium. Reproduced with permission from Petrosyan et al., Phys. Rev. Lett. 107(21), 213601 (2011). Copyright 2011 American Physical Society.Petrosyan, Otterbach, and Fleischhauer (2011)
Figure 21: (a) Rydberg-blockade-mediated interaction between slow photons. (b) Transmission versus probe detuning at various incoming photon rates. (c) Data points show photon–photon correlation function $g^{(2)}(\tau)$ at EIT resonance. Reproduced with permission from Peyronel et al., Nature 488(7409), 57-60 (2012). Copyright 2012 Springer Nature.Peyronel et al. (2012)
V.1 Nonlocal nonlinear optics
Typical EIT experiments are performed with atoms in low-lying states driven into the three-level $\Lambda$ configuration involving an intermediate short-lived excited state and two long-lived ground states.Harris, Field, and Imamoğlu (1990); Boller, Imamoğlu, and Harris (1991) When a ground state is replaced with a Rydberg state, it is viable to realize an intriguing transfer from the linear EIT in the $\Lambda$ configuration to the nonlinear EIT in the ladder configuration, as a result of both long-lifetime and strong-interaction features of highly excited Rydberg atoms.
In 2010, Adams’ group observed in 87Rb atoms a cooperative optical nonlinearity arising from Rydberg interactions and manifesting itself as a gradual reduction of the probe transmission in the EIT window when the probe intensity increasesPritchard et al. (2010) as shown in Fig. 19. The study was performed with an EIT ladder system involving ground $|g\rangle$, intermediate $|e\rangle$, and upper $|r\rangle$ levels and was explained by a simple two-atom model where one atom contributes to the three-level EIT system while the other acts as a two-level absorbing system due to dipole blockade. This fascinating experiment soon was better explained by a mean-field superatom model where the collective states of $n_{SA}=\frac{4}{3}\pi\rho R_{SA}^{3}$ atoms within a blockade sphere of radius $R_{SA}=(C_{6}\gamma_{e}/\Omega_{c}^{2})^{1/6}$ defined as the interatomic distance at which the vdWs interaction equals the half-width of steady-state population of a single atomic Rydberg state,Qian et al. (2013) are considered but appropriately truncatedPetrosyan, Otterbach, and Fleischhauer (2011) as shown in Figs. 20(a) and 20(b). This truncation is done based on two facts: i.e., (i) at most one atom is allowed to be in the upper Rydberg state and (ii) the excitation to the intermediate state $|e\rangle$ is well suppressed due to quantum destructive interference. Then it is appropriate to consider only the three lowest-order collective states $|G\rangle=|g_{1},g_{2},...,g_{n_{SA}}\rangle$, $|E^{(1)}\rangle=\frac{1}{\sqrt{n_{SA}}}\sum_{j}^{n_{SA}}|g_{1},...,e_{j},...g_{n_{SA}}\rangle$, and $|R^{(1)}\rangle=\frac{1}{\sqrt{n_{SA}}}\sum_{j}^{n_{SA}}|g_{1},...,r_{j},...g_{n_{SA}}\rangle$ for each superatom as one attempts to estimate the Rydberg excitation probability $\Sigma_{RR}$. With $\Sigma_{RR}$, it is easy to calculate the conditional polarizability $\alpha=\Sigma_{RR}\alpha_{TLA}+(1-\Sigma_{RR})\alpha_{EIT}$ with $\alpha_{TLA}$ and $\alpha_{EIT}$ being the polarizabilities for the two-level absorbing and ladder EIT systems, jointly determining the probe field propagation in the Rydberg medium. Such a computationally efficient model is in quantitative agreement with the experimental results and also predicts in Fig. 20(c) a remarkable antibunching effect in terms of the intensity correlation function $g^{(2)}_{p}(L)$ obtained at the exit from the medium, accompanied by a transmission reduction due to cooperative optical nonlinearity.
Through the utilization of Rydberg EIT, Vuletić and his colleagues showcased their ability to manipulate light fields at the level of single quanta. In their work Ref. Peyronel et al., 2012, strongly interacting Rydberg states play a crucial role in generating quantum nonlinearity via the photon-photon blockade mechanism, as shown in Fig. 21. This blockade prevents the transmission of multi-photon states by precluding the simultaneous excitation of two Rydberg atoms within the blockade radius $r_{b}$. When an incident single photon is converted into a Rydberg polariton inside the medium under EIT conditions, the Rydberg blockade prevents a second polariton from traveling within the blockade radius of the first one, leading to the destruction of EIT. The interaction between Rydberg atoms creates a strong nonlinearity that results in the attenuation of a second photon approaching the single Rydberg polariton. This phenomenon is based on the physical principle that when the blockade radius exceeds the resonant attenuation length of the medium in the absence of EIT, two photons in a tightly focused beam cannot pass through each other or propagate close to each other inside the medium. By utilizing Rydberg states of $|100S_{1/2}\rangle$, the researchers were able to realize blockade radii $r_{b}=13~{}\mu\rm m$, effectively creating a quantum nonlinear absorption filter that converts incident coherent field into a train of non-classical, anti-bunched light pulses. The demonstration of quantum-by-quantum control over light fields in this research opens up new possibilities for non-linear quantum optics based on interacting Rydberg atoms.
Soon afterward, higher-order collective states were included in an improved superatom model by going beyond the weak-probe approximation, which just results in quantitative differences, especially as the probe and coupling fields are off two-photon resonance, yet without changing main conclusions.Liu et al. (2014) Such a superatom model has been further extended to realize the polarization-selective nonlinearity and cooperative nonlinear grating unattainable in normal atoms.Wu, Artoni, and La Rocca (2015); Liu et al. (2016); Asghar et al. (2016) A semianalytical rate-equation model has also been developed for revealing a universal relation between the nonlinear optical susceptibility and the Rydberg excitation density as interatomic coherence is well suppressed due to strong dephasing effects or very different Rabi frequencies of the probe and coupling fields.Gärttner et al. (2014) Moreover, we note that a many-body iterative approach has been developed to account for the interparticle Rydberg correlations, though it is applicable only for dilute samples with less than one atom inside a superatom so that Rydberg interactions between different superatoms become dominant to reduce the many-body computational complexity. This alternative approach allows one to separate a local linear susceptibility and a nonlocal nonlinear susceptibility comparable in magnitude, hence achieving huge nonlocal optical nonlinearities beneficial, e.g., to largely enhance the Goos-Hanchen lateral shift and correlated biphoton generation.Bai and Huang (2016); Bai, Li, and Huang (2019); Zheng et al. (2022); Zhao et al. (2023)
Figure 22: (a) A light storage and retrieval experiment based on the nonlinear Rydberg EIT technique. (b) Normalized and background subtracted counts of an input probe pulse when slowly propagating or stored for $t_{s}=600$ ns. (c) Output photon number normalized by linear process efficiency $T$ against input photon number for the slow-light case and two storage times. Straight lines represent the linear behavior (oblique) and the saturation level (horizontal). Reproduced with permission from Distante et al., Phys. Rev. Lett. 117(11), 113001 (2016). Copyright 2016 American Physical Society.Distante et al. (2016)
An important application of the linear EIT in normal atoms is to realize the dynamically reversible storage and retrieval of a (classical or quantum) probe pulse by adiabatically switching off and on a control field. Such an important coherent control of the slow-light propagation and storage dynamics has been extended to Rydberg atoms exhibiting nonlocal nonlinear EIT responses. Many experiments show that storing an input probe pulse as a collective Rydberg excitation strongly enhances the Rydberg-mediated interaction compared to that in the slow-light dark-state polariton case. For example, Distante et al. observed a strong enhancement of interatomic interactions due to the largely reduced polariton group velocity for short storage times while a weak enhancement dominated by Rydberg-induced dephasing of the polariton’s multiparticle components for longer storage times.Distante et al. (2016) This conclusion is supported by Fig. 22 where the output photon number $N_{\rm out}$ depends on the input photon number $N_{\rm in}$ clearly in a nonlinear way and becomes saturated for a sufficiently large $N_{\rm in}$ as described by the input-output relation $N_{\rm out}=N_{\rm max}T(1-e^{-N_{\rm in}/N_{\rm max}})$ with $T$ being the linear process efficiency for small $N_{\rm in}$. This nonlinear light storage experiment can be well explained by extending the superatom model from its steady-state solutions to the dynamic evolution case Tian et al. (2018) which further predicts an inhomogeneous nonclassical antibunching feature of the retrieved light signal. Another effective method has also been developed to reveal the correlated photon dynamics in dissipative Rydberg media valid in the limit of a large optical depth per blockade radius.Gorshkov, Nath, and Pohl (2013); Zeuthen et al. (2017)
Figure 23: Experimental realization of contactless nonlinear optics mediated by long-range Rydberg interactions. Reproduced with permission from Busche et al., Nature Physics 13(7), 655-658 (2017). Copyright 2017 Springer Nature.Busche et al. (2017)
Even in the absence of a common medium, creating effective photon-photon interactions remains feasible.Yang et al. (2016) Busche et al.’s study delved into a novel approach to such photon interactions utilizing Rydberg interactions.Busche et al. (2017) As depicted in Fig. 23, photons stored as collective Rydberg excitations in spatially separated optical media interact through vdw interactions, inducing non-uniform phase shifts that alter the retrieval modes of the photons. These interaction-induced phase shifts are pivotal in shaping the overall interaction between the photons, resulting in observable effects like reduced retrieval in the photons’ initial modes and spatial anti-correlations in observed photon statistics. The experimental arrangement utilized highly focused signal beams and tiny optical tweezers to confine cold atoms, demonstrating the high level of precision and control in investigating non-contact nonlinear optics. Results exhibit strong agreement between theory and experiment across varying interaction strengths and distances, indicating the feasibility of implementing contactless effective photon-photon interactions in practical photonic devices. This contactless characteristic of interactions unveils new opportunities for scalable multichannel photonic devices, exploration of strongly correlated many-body dynamics using light, and the development of advanced quantum technologies based on photon interactions mediated by Rydberg states.
V.2 Single-photon engineering
Rigid dipole blockade effect in Rydberg atoms allows one to attain efficient single-photon engineering, including single-photon generation and storage due to collectively enhanced strong couplings between single Rydberg superatoms and few-photon light pulses.Paris-Mandoki et al. (2017) In Ref. Saffman and Walker, 2002, Saffman and Walker conducted a detailed analysis of collective emission and proposed the protocol for realizing deterministic single-photon source using Rydberg superatom. Instead of using Rydberg EIT to store a single photon, they show that the Rydberg superatom can be coherently prepared by using two-photon excitation to drive a many-body Rabi oscillation, hence efficiently transferred into a highly directional single photon. Honer et al.Honer et al. (2011) and Tresp et al.Tresp et al. (2016) further examined photon interactions with a Rydberg superatom under controlled dephasing conditions and discovered that induced dephasing may dictate the absorption of a single photon from any probing field, which was recently employed in the controlled multi-photon subtraction process.Stiesdal et al. (2021)
Figure 24: Single-photon source with Rydberg superatom. (a) Illustration of experimental protocol. (b) Relevant atomic levels. (c) Measured second-order-correlation function $g^{2}(0)$ as a function of the effective principal quantum number. Reproduced with permission from Dudin et al., Science 336(6083), 887–889 (2012). Copyright 2012 The American Association for the Advancement of Science.Dudin and Kuzmich (2012)
A deterministic single-photon source based on Rydberg superatom was experimentally realized in 2012. Dudin and Kuzmich (2012) As shown in Fig. 24, Dudin and Kuzmich performed 780 nm-480 nm two-photon Rydberg excitation and photon retrieval in an ensemble of 87Rb atoms confined within the blockade radius. When high-lying Rydberg states ($n\sim 100$) are used, the measured second-order-correlation function $g^{2}(0)$ of the retrieved photon approaches zero, indicating the successful preparation of a single photon. Further research in Rydberg single-photon source focused on analyzing and improving the quality of generated single photons.Lampen et al. (2019); Ornelas-Huerta et al. (2020); Padrón-Brito et al. (2021); Shi et al. (2022) The state-of-the-art Rydberg single-photon sources feature near-unity purity and indistinguishability, which are comparable to those in other leading single-photon physical platformsAharonovich, Englund, and Toth (2016) and have been used to achieve a high-fidelity photonic quantum logic gate using linear optical approach.Shi et al. (2022)
Figure 25: Transition $|u\rangle\rightarrow|d\rangle$ of a source atom (top left) is coupled nonresonantly to transition $|i\rangle\rightarrow|s\rangle$ of the medium atoms (top right) via DDE interaction $D$. A laser pulse $\Omega$ couples ground state $|g\rangle$ to Rydberg state $|i\rangle$ and leads to the transition $|g\rangle\rightarrow|s\rangle$ with the help of DDE transition $|i\rangle|u\rangle\rightarrow|s\rangle|d\rangle$. The resulting single Rydberg excitation exhibits the profile $S(\textbf{r})\propto D(\textbf{r}-\textbf{r}_{s})$ and can be converted into a propagating photon $\mathcal{E}$ by applying the control field $\Omega_{c}$. Reproduced with permission from Petrosyan et al., Phys. Rev. Lett. 121(12), 123605 (2018). Copyright 2018 American Physical Society.Petrosyan and Mølmer (2018)
Figure 26: (a) Level scheme, (b) simplified schematic, and (c) pulse sequence for the realization of an all-optical transistor with $U(r)$ denoting the dipole-dipole interaction between two atoms of distance $r$ and in states $|r_{g}\rangle$ and $|r_{s}\rangle$, respectively. (d) Absorption spectrum of the source field (dots) exhibiting an EIT window at $\delta_{s}=0$ and that of the gate field (circles) around $\delta_{g}=40$ MHz. The contrast of the switches (red solid) against mean numbers of (e) incident and (f) stored gate photons in coherent states is presented with fundamental limits (pink dashed) set by the corresponding photon statistics. (g) Gate photon optical gain and (h) source photon transfer function with $\overline{N}_{\rm g,in}=0$ (red) and $\overline{N}_{\rm g,in}=0.75$ (blue) for coherent states, while $N_{\rm in,Fock}=1$ (green) and $N_{\rm st,Fock}=1$ (black) for Fock states. Reproduced with permission from Gorniaczyk et al., Phys. Rev. Lett. 113(5), 053601 (2014). Copyright 2014 American Physical Society.Gorniaczyk et al. (2014)
In 2018, Petrosyan and Mølmer proposed a more efficient free space scheme to create deterministic single photons in a well-defined spatio-temporal mode.Petrosyan and Mølmer (2018) This is done by first excite a single source atom into a high-lying Rydberg state $|u\rangle$, which couples to an ensemble of ground state atoms via dipole-dipole exchange (DDE) interactions so that its transition to another Rydberg state $|d\rangle$ is accompanied by the creation of a spatially extended spin wave $S(\textbf{r})$ of a single Rydberg excitation in the ensemble, as shown in Fig. 25. This collective single-Rydberg excitation can be converted to a propagating optical photon $\mathcal{E}$ as a coherent control field $\Omega_{c}$ is switched on to form the ladder EIT configuration. This nontrivial scheme relies on neither the strong coupling of a single emitter to a resonant cavity nor the heralding of collective excitation or complete Rydberg blockade of multiple excitations in a mesoscopic atomic ensemble. Similarly, it is feasible to achieve reliable quantum memory at the single-photon level by absorbing all extra photons inside each blockade volume of the Rydberg superatoms.Li et al. (2014); Li and Kuzmich (2016); Distante et al. (2017); Zhang et al. (2022b)
The dipole blockade and exchange mechanisms described above used for generating single-photon sources can also be extended to design usually unattainable single-photon devices, including single-photon switch,Baur et al. (2014) single-photon transistor, Gorniaczyk et al. (2014); Tiarks et al. (2014) and single-photon absorber. Tresp et al. (2016); Honer et al. (2011) For instance, one may store a gate light pulse containing only one photon on average as a collective Rydberg excitation in cold atoms, in the presence of which the transmission of the subsequent target pulse is largely suppressed due to dipole blockade.Baur et al. (2014) Such a single-photon switch offers many interesting prospects ranging from quantum communication to quantum computing, such as heralding a successful quantum memory event, nondestructively detecting an optical photon, and building a photonic quantum-logic gate. Moreover, it is viable to realize an all-optical single-photon transistor by mapping gate and source (target) photons to strongly interacting Rydberg excitations with different principal quantum numbers Gorniaczyk et al. (2014) or by suppressing and enhancing the transmission of target (source) photons through dipole blockade and Förster resonances, respectively.Tiarks et al. (2014)
To be more specific, we can see from Figs. 26(a)-26(d) that an all-optical transistor can be achieved with the following two procedures. First, store photons in a gate field $\mathcal{E}_{g}$ as collective Rydberg excitation in a sample of cold 87Rb atoms by coupling the ground state $|g\rangle$ with the Rydberg state $|r_{g}\rangle$ through a control field $\Omega_{g}$. Second, a source field $\mathcal{E}_{s}$ is sent through the same medium at a reduced group velocity due to the resonant EIT provided by the control field $\Omega_{s}$ coupling to the Rydberg state $|r_{s}\rangle$. Then individual source photons can travel with little absorption through the atomic sample in the absence of a gate excitation for a low source photon rate or are strongly attenuated by the scattering from the intermediate state $|e\rangle$ in the blockade regime where Rydberg interactions between states $|r_{g}\rangle$ and $|r_{s}\rangle$ result in a two-level absorbing system. This realizes a transistor working at the single-photon level with its main performance illustrated in Figs. 26(e)-26(h). First, we can see from Fig. 26(e) that the switch contrast $C$ plotted against the mean number of incident gate photons $\overline{N}_{g,{\rm in}}$ is clearly below the fundamental limit set by the coherent-state photon statistics. Figure 26(f) shows, however, that $C$ plotted against $\overline{N}_{g,{\rm st}}$ (the mean number of stored gate photons) gets much closer to the corresponding fundamental limit. It is more important to note that, on average, about $3$ input or $2$ stored gate photons are already enough to achieve a switch contrast over $0.8$. Figure 26(g) further shows that the optical gain is up to $20\sim 30$ (limited by the self-blocking effect of the source photons) where a coherent input with $\overline{N}_{g,{\rm in}}=0.75(3)$ gate photons can be used to effectively control the transmission of $30\sim 50$ source photons. Finally, we find in Fig. 26(h) a nonlinear saturation behavior due to self-blockade in terms of the source photon transfer function from $\overline{N}_{s,{\rm in}}$ to $\overline{N}_{s,{\rm out}}$. Hence a much higher optical gain is expected to occur provided the source photon loss due to self-blockade can be overcome.
[…]
V.3 Hybrid Rydberg systems
Hybrid quantum systems are systems that integrate various types of quantum components or platforms to capitalize on their complimentary qualities. These systems seek to exploit the distinct qualities of diverse quantum technologies in order to produce more robust, scalable, and adaptable quantum devices. Hybrid systems involving Rydberg atoms promise intriguing possibilities for quantum information processing. It was shown that a JC model could be realized in the optical domain for an ensemble of atoms by exploiting their collective couplings to a quantized cavity mode.Guerlin et al. (2010); Shao, Wu, and Yi (2017) This is realized by restricting the huge state space to the ground state $|G\rangle$ and the collective excited state $|R\rangle$ with a single Rydberg excitation due to Rydberg blockade. The main benefit is that the collectively enhanced coupling between the emitter and the cavity mode can be much larger than the decay rates, leading to the strong coupling regime of an effective JC model. Trapping some atoms in a volume smaller than the blockade sphere close to an atom chip, it is viable to attain a hybrid system consisting of a superatom coupled to a surface phonon polariton, which can be used to subtract individual photons from a beam of light.Kübler et al. (2013) Hybrid systems consisting of macroscopic mechanical oscillators coupled to a small ensemble of Rydberg atoms are also widely investigated to achieve ground state cooling, quantum state engineering, nontrivial nonlinear dynamics, bistable Rydberg excitation, and mechanical oscillation, etc.Yan et al. (2015); Bariani et al. (2014); Carmele et al. (2014) This depends especially on the collectively enhanced coupling arising from the dipole blockade effect of Rydberg superatoms.Stiesdal et al. (2018); Yang et al. (2022b) The strong coupling regime may also be realized between a Rydberg atomic ensemble and propagating surface phonon polaritons on a piezoelectric superlattice.Sheng, Chao, and Shaffer (2016) These hybrid Rydberg systems provide a promising platform for extending the contents of nonlinear quantum optics and photonics, essential for developing future quantum control techniques. Moreover, when combined with Floquet engineering techniques, polaritons, formed by photons entering a cavity and hybridized with atomic Rydberg excitations, can be efficiently manipulated to interact with each other.Clark et al. (2019) They can also be simulated to exhibit topological order, such as the Laughlin state.Clark et al. (2020)
Figure 27: The experimental setup involves measuring transmission, with EIT spectra depicted in blue (presence of EIT control) or red (absence of EIT control), normalized against the maximum transmission of an empty cavity (depicted in gray). Reproduced with permission from Vaneecloo et al., Phys. Rev. X 12(2), 021034 (2022). Copyright 2022 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Vaneecloo, Garcia, and Ourjoumtsev (2022)
Figure 28: (a) A bow-tie cavity composed of two concave mirrors and two convex mirrors. Relative intensity (b) and (c) phase of signal light as a function of signal detuning referring to cavity-enhanced Rydberg EIT spectra. Red (blue) data and lines are attained in the absence (presence) of coupling light. Reproduced with permission from Stolz et al., Phys. Rev. X 12(2), 021035 (2022). Copyright 2022 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Stolz et al. (2022)
Below, we will explain the above discussions in more detail by focusing on recent experiments.Vaneecloo, Garcia, and Ourjoumtsev (2022); Stolz et al. (2022); Magro et al. (2023) As can be seen in Fig. 27, by placing a small cloud of 87Rb atoms inside a cavity with a finesse of $\mathcal{F}=590$, one observes the vacuum Rabi splitting of two transmission peaks located at $\delta/2\pi=\pm 10$ MHz, indicating a cavity-mode coupling constant ($g/2\pi=10$ MHz) exceeding both the cavity decay rate ($\kappa/2\pi=2.9$ MHz) and the atomic decay rate ($\gamma/2\pi=3$ MHz). Exciting these atoms to a high Rydberg state $|R^{\prime}\rangle$ by the EIT control and D1 probe beams, a narrower EIT window is found to open on resonance corresponding to a dark polariton state mixing a photon in the cavity and a Rydberg excitation in the atomic cloud. On the other hand, the $109S$ drive and D2 drive beams can drive these atoms to another high Rydberg state $|R\rangle$ via a two-photon Raman resonance so that they behave as a two-level superatom due to the blockade of multiple Rydberg excitations. Atoms in states $|R\rangle$ and $|R^{\prime}\rangle$ further strongly interact with another pair of Rydberg states through Förster resonance. Then, by shifting the state $|R\rangle$ out of resonance, a superatom in the state $|R^{\prime}\rangle$ destroys the EIT response to the D1 probe, which drives the photonic component of the $|R^{\prime}\rangle$ polariton. This scheme has been exploited to realize the high-efficiency single shot optical detection of a Rydberg superatom, the superatom state-dependent $\pi$ phase shift of light reflected by the cavity, and the deterministic preparation of freely propagating photonic qubits with negative Wigner functions,Vaneecloo, Garcia, and Ourjoumtsev (2022); Magro et al. (2023) by taking advantage of two complementary approaches: i.e., cavity quantum electrodynamics and interacting atomic ensembles.Kumar et al. (2016)
It is also worth noting that a bow-tie cavity has been incorporated into a Rydberg EIT experiment for demonstrating a polarization CNOT gate between two optical pulses interacting with hundreds of trapped 87Rb atoms at an average efficiency of about $41.7\%$, far exceeding that in previous experiments.Stolz et al. (2022) This atom-cavity system in a ring geometry minimizes dephasing resulting from the combination photon recoil of atomic motion and is characterized by $(g,\kappa,\gamma)=2\pi\times(1.0,2.3,3.0)$ MHz, thus yielding finesse $\mathcal{F}=350$ and a collective cooperativity of $21$ facilitating the enhanced interaction of single photons and Rydberg atoms. The four-mirror bow-tie cavity is shown in Fig. 28(a) where the target and control coupling beams are guided into the cavity from different concave mirrors, while the target and control qubits are input from a common convex mirror. This ring resonator is necessary for converting the $\pi$ conditional phase shift into a two-photon gate for polarization qubits. The relative intensity and phase of the target signal reflected from the cavity have been measured as a function of the signal detuning from the cavity resonance in Figs. 28(b) and 28(c), respectively. In the absence of an EIT coupling light, the intensity exhibits a normal-mode splitting, while the phase shows two dispersive features, one at a normal mode (see red data). As the EIT coupling light is on, a narrow EIT feature appears at resonance so that a $2\pi$ phase boost may occur as the signal detuning is tuned through zero (see blue data). This forms the basis for a $\pi$ conditional phase shift, essential for constructing a CNOT gate.
Another interesting hybrid setup involves polar molecules and Rydberg atoms trapped in optical tweezer arrays in which the Rydberg blockade due to the charge-dipole interaction is demonstrated. Guttridge et al. (2023) This setup is highly beneficial because quantum information can be encoded in the internal states of the molecule, and quantum gates can be performed by utilizing the strong Rydberg-Rydberg interactions of atoms, with little sensitivity to the motional states of the particles. Kuznetsova et al. (2011); Wang et al. (2022); Zhang and Tarbutt (2022)
VI CONCLUSION AND OUTLOOK
In conclusion, Rydberg atoms have emerged as one of the most promising platforms for probing quantum physics and quantum technological applications. The essential elements necessary for understanding Rydberg physics are covered in the first part of this review article, particularly the optical excitation of Rydberg states and the emergence of different types of Rydberg-Rydberg interactions, such as vdw and dipolar interactions, including the tunability using external fields. The strong interatomic interactions lead to the emergence of superatoms in Rydberg gases via the Rydberg blockade. When driving by resonant lasers, many-body Rabi oscillations can be found. The superatom approach is an efficient way to describe the dynamics and collective states, while the underlying mechanism can be used to generate robust and scalable entanglement in neutral atoms. Further, the same mechanism can create the entanglement between atoms and photons and between two photons. In an atom array, Rydberg blockade and superatoms serve a mechanism to emulate many-body spin models to create various many-body states and intricate quantum dynamics.
The remarkable collective excitation properties of Rydberg atoms hold immense promise for advancing quantum information processing and quantum optics. It is expected that an increasing number of experimental achievements and theoretical proposals will be seen in the coming years. Future developments have a number of exciting perspectives. Quantum simulation, quantum computation and quantum optics applications are promising growing area. To fully unlock the potential of Rydberg atom systems, new schemes and experimental techniques are needed in order to mitigate limitations (such as spin wave dephasing) in existing experiments. Moreover, by combining different quantum systems, their individual properties can be leveraged in order to realize unique functionality. Recently, hybrid systems in which Rydberg excitations combine with Bose-Einstein condensates, Heidemann et al. (2008); Rammohan et al. (2021) trapped ions, Zhang et al. (2020b); Mokhberi, Hennrich, and Schmidt-Kaler (2020) polar molecules, Guttridge et al. (2023); Zhang and Tarbutt (2022); Wang et al. (2022) mechanical oscillators, Antezza et al. (2014) etc., pave the way for the exploration of novel applications and physics in the future.
ACKNOWLEDGMENTS
The authors extend their sincere gratitude to the anonymous reviewers for their thorough reading and valuable suggestions, which have significantly improved the readability of the manuscript. Additionally, we would like to thank F. Cesa, C. S. Adams, S. Dürr, S. Hofferberth, Y. O. Dudin, D. Petrosyan, J. Zeiher, and A. Ourjoumtsev for their valuable comments. This work was supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12174048, 12274376, 12074061, 12374329, and U21A6006, and the National Key Research and Development Program of China under Grant No. 2021YFA1402003. R.N. acknowledges DST-SERB for Swarnajayanti fellowship File No. SB/SJF/2020-21/19, the MATRICS grant (MTR/2022/000454) from SERB, and funding from National Mission on Interdisciplinary Cyber-Physical Systems (NM-ICPS) of the Department of Science and Technology, Govt. of India through the I-HUB Quantum Technology Foundation, Pune INDIA. W.L. acknowledges support from the EPSRC through Grant No. EP/W015641/1 and the Going Global Partnerships Programme of the British Council (Contract No. IND/CONT/G/22-23/26).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Xiao-Qiang Shao: Project administration (equal); Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Shi-Lei Su: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Lin Li: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Rejish Nath: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Jin-Hui Wu: Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal). Weibin Li: Project administration (equal); Funding acquisition (equal); Writing-original draft (equal); Writing-review & editing (equal).
DATA AVAILABILITY
The data of Figs. 1 and 5 generated in this review article are available from the corresponding authors upon reasonable request.